### v1

# Linear Bandits with Partially Observable Features

Abstract
We introduce a novel linear bandit problem with partially observable features, resulting in partial reward information and spurious estimates. Without proper address for latent part, regret possibly grows linearly in decision horizon $T$, as their influence on rewards are unknown. To tackle this, we propose a novel analysis to handle the latent features and an algorithm that achieves sublinear regret. The core of our algorithm involves (i) augmenting basis vectors orthogonal to the observed feature space, and (ii) introducing an efficient doubly robust estimator. Our approach achieves a regret bound of $\widetilde{O}(\sqrt{(d+d_{h})T})$, where $d$ is the dimension of observed features, and $d_{h}$ is the unknown dimension of the subspace of the unobserved features. Notably, our algorithm requires no prior knowledge of the unobserved feature space, which may expand as more features become hidden. Numerical experiments confirm that our algorithm outperforms both non-contextual multi-armed bandits and linear bandit algorithms depending solely on observed features.
Keywords: linear bandits, partially observable features, doubly robust estimation
1 Introduction
We consider a linear bandit problem where the learning agent has access to only a subset of the features, while the reward is determined using the complete set of features, including both observed and unobserved elements  [^1] In this paper, we use the terms “unobserved” and “latent” interchangeably to refer to features that are not visible to the decision-making agent.. Conventional linear bandit problems rely on the assumption that the rewards are linear to only observed features, without accounting for the potential presence of unobserved features. However, in many real-world applications, rewards are often affected by the latent features that are not observable to the agent. For example, in recommendation systems, the true reward — such as user satisfaction or purchase decisions — depends not only on observed features like user demographics or past behaviors but also on latent preferences, such as specific tastes in artists (for streaming services) or brands (in e-commerce). Accurately incorporating these latent features is essential for providing precise recommendations, while ignoring them causes bias or model misspecification errors in every decision-making.
To address the latent features, Park and Faradonbeh (2022), Kim et al. (2023a) and Park and Faradonbeh (2024) rely on the assumption that observed features are linear to the latent features sampled from a specific distribution, e.g., a mean-zero Gaussian. Establishing a regret bound sublinear in the decision horizon without such structural assumptions on the latent features remains a significant challenge and has not been accomplished yet. Key challenges in the bandit problem with partially observable features arise from the complete lack of information on the latent features. Indeed, we do not even know whether an agent observes features partially nor whether we should use the latent features. To address these challenges, we propose a novel linear bandit algorithm that is agnostic to partially observability. Despite the absence of knowledge regarding unobserved features, our algorithm attains a tighter regret bound than both linear bandit algorithms that considers only observed features and MAB algorithms that ignore features entirely. Specifically, our proposed algorithm achieves a $\sqrt{T}$-rate regret bound, without requiring any prior knowledge of the unobserved features, where $T$ is the decision horizon.
The key idea of our proposed algorithm lies in two main components: (i) reconstructing feature vectors to capture the influence of unobserved features on rewards, and (ii) constructing a novel doubly robust estimator to mitigate information loss due to unobservability. For (i), we decompose reward into two additive terms: one projected onto the row space of observed features, and the other onto its orthogonal complement. The former term maximally captures the effect of observed features, while the latter minimizes the impact of unobserved features. We then augment observed features with an orthogonal basis from the complement space. This allows us to reformulate the problem within a conventional linear bandit framework, where the reward function is defined as a dot product of minimally augmented features and an unknown parameter. However, since these augmented features are not identical to the unobserved features, potential estimation error may arise. To mitigate these errors, we leverage (ii) a doubly robust estimator, which is widely used in statistical literature for its robustness to errors due to missing data. These two approaches allow our algorithm to effectively compensate for missing information, improving both estimation accuracy and adaptability to the environment.
Table 1: An overview of regret bound range of our algorithm, RoLF, depending on $d_{h}\in[0,K-d]$, the dimension of the vector space spanned by the rows of the matrix of unobserved features influencing the reward. Note that $\widetilde{O}$ denotes the big-O notation omitting logarithm factors.
Feature space Regret bound
span(observed features) $\supseteq$ span(latent features) $\widetilde{O}(\sqrt{dT})$
span(observed features) $\subseteq$ span(latent features) $\widetilde{O}(\sqrt{KT})$
otherwise $\widetilde{O}(\sqrt{(d+d_{h})T})$
Our main contributions are summarized as follows:
  • We propose a linear bandit problem with partially observable features. Our problem setting is more general and challenging than those in the existing literature on linear bandits with latent features, which often rely on specific structural assumptions governing the relationship between observed and latent features. In contrast, our approach assumes no additional structure for the unobserved features beyond the linearity of the reward function, which is commonly adopted in the linear bandit literature (Section 3).
  • We introduce a novel estimation strategy by (i) efficiently augmenting the features that maximally captures the effect of reward projected onto the observed features, while minimizing the impact of unobserved features (Section 4), and (ii) constructing a doubly robust (DR) estimator that mitigates errors from unobserved features. By integrating augmented features with the DR estimator, we guarantee a $t^{-1/2}$ convergence rate on the rewards for all arms in each round $t$ (Theorem 2).
  • We propose an algorithm named Robust to Latent Features (RoLF) for general linear bandit framework with latent features (Algorithm 1). The algorithm achieves a regret bound of $\widetilde{O}(\sqrt{(d+d_{h})T})$ (Theorem 3), where $d_{h}$ is the dimension of the subspace formed by projecting the reward from unobserved features onto the orthogonal complement of the row space of observed features (Section 4.2) and $\widetilde{O}(\cdot)$ is the Big-O notation omitting logarithmic factors. RoLF requires no prior knowledge or modeling of unobserved features yet achieves a sharper regret bound than both linear bandit algorithms that consider only observed features (Li et al., 2010; Abbasi-yadkori et al., 2011; Agrawal and Goyal, 2013; Kim and Paik, 2019) and MAB algorithms (Auer et al., 2002), as summarized in Table 1.
  • Our experiments confirm that our algorithm consistently outperforms MAB and linear bandit algorithms that solely depend on observed features, validating both its practicality and theoretical guarantees.
2 Related Works
In bandit problems, the learning agent learns only from the outcomes of chosen actions, leaving unchosen alternatives unknown (Robbins, 1952). This constraint requires a balance between exploring new actions and exploiting actions learned to be good, known as the exploration-exploitation tradeoff. Efficiently managing this tradeoff is crucial for guiding the agent towards the optimal policy. To address this, algorithms based on optimism in the face of uncertainty (OFU) (Lai and Robbins, 1985) are widely used and studied in linear bandits (Abe and Long, 1999; Auer, 2002; Dani et al., 2008; Rusmevichientong and Tsitsiklis, 2010). Notable examples include LinUCB (Li et al., 2010; Chu et al., 2011) and OFUL (Abbasi-yadkori et al., 2011), known for their practicality and performance guarantees. However, existing approaches differ from ours in two key aspects: (i) they assume that the learning agent can observe the entire feature vector related to the reward, and (ii) their algorithms have regret that scales linearly with the dimension of the observed feature vector, i.e., $\widetilde{O}(d\sqrt{T})$.
In contrast, we develop an algorithm that achieves a sublinear regret bound by employing the doubly robust (DR) technique, thereby avoiding the linear dependence on the dimension of the feature vectors. The DR estimation in the framework of linear contextual bandits is first introduced by Kim and Paik (2019) and Dimakopoulou et al. (2019), and subsequent studies improve the regret bound in this problem setting by a factor of $\sqrt{d}$ (Kim et al., 2021; 2023b). A recent application (Kim et al., 2023c) achieves a regret bound of order $O(\sqrt{dT\log T})$ under IID features over rounds. However, the extension to non-stochastic or non-IID features remains an open question. To address this issue, we develop a novel analysis that applies the DR estimation to non-stochastic features, achieving a regret bound sublinear with respect to the dimension of the augmented feature vectors. Furthermore, we extend DR estimation to handle sparse parameters, thereby further improving the regret bound to be sublinear with respect to the reduced dimension.
Our problem is more general and challenging than misspecified linear bandits, where the assumed reward model fails to accurately reflect the true reward, such as when the true reward function is non-linear (Lattimore and Szepesvári, 2020), or a deviation term is added to the reward model (Ghosh et al., 2017; Bogunovic et al., 2021; He et al., 2022). While our work assumes that the misspecified (or inaccessible) portion of the reward is linearly related to certain unobserved features, misspecified linear bandit problems can be reformulated as a special case of our framework. While the regret bounds in Lattimore and Szepesvári (2020), Bogunovic et al. (2021) and He et al. (2022) incorporate the sum of misspecification errors that may accumulate over the decision horizon, our work establishes a regret bound that is sublinear in the decision horizon $T$, not affected by misspecification errors. Ghosh et al. (2017) proposed a hypothesis test to decide between using linear bandits or MAB, demonstrating an $O(K\sqrt{T}\log T)$ regret bound when the total misspecification error exceeds $\Omega(d\sqrt{T})$. In contrast, our algorithm achieves an $O(\sqrt{(d+d_{h})T\log T})$ regret bound without requiring hypothesis tests for misspecification or partial observability.
Lastly, our problem appears similar to the bandits with partially observable features, as studied by Park and Faradonbeh (2022) and Zeng et al. (2024). Both works assume that observed features are related to latent features through a known linear mapping, with the latent features sampled from a certain Gaussian distribution. Notably, the latter work further assumes that the latent features evolves following a specific linear dynamics model. In addition, both approaches aim to recover the latent features: the former introduces a known decoder mapping from observed features to their corresponding latent features, while the latter estimates them using a Kalman filter. In contrast, our approach imposes no structural assumptions on either the observed or latent features, making our problem more general and fundamentally more challenging than those addressed in these works. Furthermore, our work does not attempt to recover any information related to latent features. Instead, our novel approach demonstrates that the best action can still be identified and chosen, even with an unobserved portion of the reward, by solely exploiting observed features.
3 Preliminaries
3.1 Notation
For any $n\in\mathbb{N}$, let $[n]$ denote the set $\{1,2,\dots,n\}$. For a vector $\mathbf{v}$, we denote its $L_{1}$, $L_{2}$ and supremum norm by $\|\mathbf{v}\|_{1}$, $\|\mathbf{v}\|_{2}$, and $\|\mathbf{v}\|_{\infty}$, respectively. The $L_{2}$-norm weighted by a positive definite matrix $\mathbf{D}$ is denoted by $\|\mathbf{v}\|_{\mathbf{D}}$. For two vectors $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$, the inner product is defined as the dot product, i.e., $\langle{\mathbf{v}_{1}},{\mathbf{v}_{2}}\rangle:=\mathbf{v}_{1}^{\top}\mathbf{v}_{2}$, and we use both notations interchangeably. For a matrix $\mathbf{M}$, its minimum and maximum eigenvalue are denoted by $\lambda_{\min}(\mathbf{M})$ and $\lambda_{\max}(\mathbf{M})$, respectively. We let $\mathrm{R}(\mathbf{M})$ denote the row space of $\mathbf{M}$, i.e., a subspace spanned by the rows of $\mathbf{M}$.
3.2 Problem Formulation
In this section, we outline our problem setting and introduce several key assumptions. Each arm $a\in[K]$ is associated with a true feature vector $\mathbf{z}_{a}\in\mathbb{R}^{d_{z}}$ that determines the rewards. However, the agent can observe only a subset of its elements, with the others remaining unobserved. Specifically, $\mathbf{z}_{a}$ is defined as follows:
$\mathbf{z}_{a}:={\left[{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}x_{a}^{(1)}},\cdots,{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}x_{a}^{(d)}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}u_{a}^{(1)}},\cdots,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}u_{a}^{(d_{u})}}\right]}^{\top},$ (1)
For clarity, we highlight the observed components in blue, while the unobserved components in red. Note that the dimensions of the latent feature vector, $d_{u}=d_{z}-d$, and the true feature vector, $d_{z}$, are both unknown to the agent. Consequently, the agent is unaware of whether the features are partially observed, which introduces significant challenges in selecting appropriate strategies.
It is worth noting that the setting with fixed observed features [^2] This assumption is standard in linear bandits with model misspecification (Ghosh et al., 2017; Lattimore et al., 2020), which is a special case of our partially observable feature setting. includes linear bandits with misspecification error (Ghosh et al., 2017; Bogunovic et al., 2021; He et al., 2022) as special cases. In Appendix B, we present a setting with varying observed features and an algorithm that achieves $\sqrt{T}$-rate regret bound. Moreover, if latent features were allowed to change arbitrarily over time, the problem would become non-learnable and thus ill-posed. Consequently, assuming fixed features is both natural and well-justified (see Table 2 for comparisons).
Table 2: Summary of problem settings covered in this paper and the corresponding results. Note that if latent features arbitrarily change over time, the problem itself would become non-learnable, making the problem ill-posed (see Appendix B for details).
Observed Features Unobserved Features Learnable? Results
Fixed Fixed Yes Theorem 3
Varying Fixed Yes Theorem 6
Varying Varying No -
The reward associated with each arm is defined as the dot product of its true features $\mathbf{z}_{a}$ and an unknown parameter $\boldsymbol{\theta}_{\star}\in\mathbb{R}^{d_{z}}$, given by $y_{a,t}=\langle{\mathbf{z}_{a}},{\boldsymbol{\theta}_{\star}}\rangle+\epsilon_{t}=\mathbf{z}_{a}^{\top}\boldsymbol{\theta}_{\star}+\epsilon_{t}$ for all $a\in[K]$. The error term, $\epsilon_{t}$, captures the inherent randomness in the reward, and we adopt a standard assumption commonly used in bandit problems:
Assumption 1 (Sub-Gaussian noise).
Let $\{\mathcal{F}_{t}\}_{t\in[T]}$ denote history at round $t$, represented by a filtration of sigma algebras. The reward noise $\epsilon_{t}$ is assumed to be a $\sigma$-sub-Gaussian random variable conditioned on $\mathcal{F}_{t}$. Formally,
$\mathbb{E}[{\exp(\lambda\epsilon_{t})}{\hskip 0.80002pt|\hskip 0.80002pt}{\mathcal{F}_{t-1}}]\leq\exp\left(\lambda^{2}\sigma^{2}\over 2\right),$
for all $\lambda\in\mathbb{R}$.
It follows that $\mathbb{E}[{\epsilon_{t}}{\hskip 0.80002pt|\hskip 0.80002pt}{\mathcal{F}_{t-1}}]=0$, and $\mathbb{E}[{y_{a,t}}{\hskip 0.80002pt|\hskip 0.80002pt}{\mathcal{F}_{t-1}}]=\langle{\mathbf{z}_{a}},{\boldsymbol{\theta}_{\star}}\rangle$ under this assumption. For brevity, we use $\mathbb{E}_{t-1}[\cdot]$ to denote $\mathbb{E}[{\cdot}{\hskip 0.80002pt|\hskip 0.80002pt}{\mathcal{F}_{t-1}}]$ henceforth. Given that $\epsilon_{t}$ is sampled after each action is observed, $\epsilon_{t}$ is $\mathcal{F}_{t}$-measurable. To eliminate issues of scale for analysis, we assume that the expected reward $|\langle\mathbf{z}_{a},\boldsymbol{\theta}_{\star}\rangle|\leq 1$ for all $a\in[K]$.
Let $a_{\star}:=\mathop{\mathrm{argmax}}_{a\in[K]}\langle{\mathbf{z}_{a}},{\boldsymbol{\theta}_{\star}}\rangle$ denote the optimal action, considering both observed and latent features. The theoretical performance of our algorithm is evaluated through cumulative regret, which measures the total expected difference between the reward of the optimal action and the reward of the action selected in each round. Formally,
$\begin{split}\mathrm{Reg}(T)&=\mathbb{E}\left[\sum_{t=1}^{T}\langle{{\mathbf{z}_{\star}}-{\mathbf{z}_{t}}},{\boldsymbol{\theta}_{\star}^{(o)}}\rangle\right]\end{split}.$ (2)
Considering the composition of $\mathbf{z}_{a}$ defined in Eq. 1, we can decompose the parameter as $\boldsymbol{\theta}_{\star}=[({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\boldsymbol{\theta}_{\star}^{(o)}})^{\top},({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\boldsymbol{\theta}_{\star}^{(u)}})^{\top}]^{\top}$, where $\boldsymbol{\theta}_{\star}^{(o)}\in\mathbb{R}^{d}$ and $\boldsymbol{\theta}_{\star}^{(u)}\in\mathbb{R}^{d_{u}}$ are the parameters for observed and latent features, respectively. Adopting this, the reward $y_{a_{t},t}$ can also be decomposed into three terms:
$y_{a_{t},t}=\langle{\mathbf{x}_{a_{t}}},{\boldsymbol{\theta}_{\star}^{(o)}}\rangle+\epsilon_{t}+\langle{\mathbf{u}_{a_{t}}},{\boldsymbol{\theta}_{\star}^{(u)}}\rangle,$ (3)
where the last term corresponds to the inaccessible portion of the reward. This reward model is equivalent to that imposed in the linear bandits with misspecification error (Lattimore et al., 2020). While the regret bound in Lattimore et al. (2020) includes misspecification error that grows linearly in decision horizon, our proposed method (Section 4) addresses this misspecification error and achieves a regret bound that is sublinear in the decision horizon.
Before presenting our method and algorithm, we first establish a lower bound for the regret incurred by algorithms that disregard the unobserved portion of rewards. Specifically, the following theorem establishes a lower bound for two algorithms relying solely on observed features: OFUL (Abbasi-yadkori et al., 2011) and LinTS (Agrawal and Goyal, 2013).
Theorem 1 (Regret lower bound of OFUL and LinTS ignoring latent features).
There exists a problem instance where the expected regret of both OFUL (LinUCB) and LinTS grows linearly in $T$.
Sketch of proof.
Consider a linear bandit problem with action set $\mathcal{A}:=\{1,2\}$, where true feature set is defined as $\mathcal{Z}:=\{[1,3]^{\top},[2,19/4]^{\top}\}\subset\mathbb{R}^{2}$ and the arm 2 is the optimal action. The latent portion of each reward is “large”, in the sense that there exists a constant $C>0$ such that $|\langle{\mathbf{u}_{a_{t}}},{\boldsymbol{\theta}_{\star}^{(u)}}\rangle|=|\mathbb{E}[y_{a_{t},t}]-\langle{\mathbf{x}_{a_{t}}},{\boldsymbol{\theta}_{\star}^{(o)}}\rangle|>C$ (following Eq. 3). Under this setup, the estimator associated with observed features is not consistent and the OFUL and LinTS selects the suboptimal arm with probability $\Theta(1)$. Consequently, the regret grows linearly in $T$ for both cases. ∎
Theorem 1 implies that neglecting the latent portion of the reward in decision-making could result in a failure in the learning process of the agent. The comprehensive proof is deferred to Section C.1. While Theorem 1 focuses on OFUL and LinTS, which are known to achieve the most efficient regret bounds for UCB and Thompson Sampling-based policies, we prove an algorithm-agnostic lower bound using different analysis (see Appendix D for details).
4 Robust Estimation for Partially Observable Features
We propose our estimation method to obtain a sublinear regret bound for linear bandits with latent features. Section 4.1 introduces the feature vector augmentation to handle the misspecification error and Section 4.2 presents the doubly robust estimation to further improve the regret bound.
4.1 Feature Vector Augmentation with Orthogonal Projection
In order to minimize regret, it is sufficient to estimate the $K$ expected rewards $\{\mathbf{z}_{a}^{\top}\boldsymbol{\theta}_{\star}:a\in[K]\}$ rather than all components of $\boldsymbol{\theta}_{\star}\in\mathbb{R}^{d_{z}}$. A straightforward approach to this problem, which achieves a regret bound of $\widetilde{O}(\sqrt{KT})$, is to disregard the observed features and apply MAB algorithms like UCB1 (Auer et al., 2002). However, these algorithms tend to incur higher regret than those leveraging features, particularly when the number of arms is significantly larger than the dimension of the feature vectors, i.e., $K\gg d$.
We propose a unified approach to handle all cases of partially observable features and efficiently estimate all $K$ expected rewards. Let $\mathbf{X}:=(\mathbf{x}_{1},\dots,\mathbf{x}_{K})\in\mathbb{R}^{d\times K}$ represent a matrix that concatenates the observed part of the true features, and $\mathbf{U}:=(\mathbf{u}_{1}^{(u)},\ldots,\mathbf{u}_{K}^{(u)})\in\mathbb{R}^{d_{u}\times K}$ represent the matrix that concatenates the latent complements of the true features for each arm. We assume a set of $K$ vectors $\{\mathbf{x}_{1},\ldots,\mathbf{x}_{K}\}$ spans $\mathbb{R}^{d}$, without loss of generality. [^3] When $d>K$, we can apply singular value decomposition on $\mathbf{X}$ to reduce the feature dimension to $\bar{d}\leq K$ with $\mathrm{R}(\mathbf{X})=\bar{d}$. We define $\mathbf{P}_{\mathbf{X}}:=\mathbf{X}^{\top}(\mathbf{X}\mathbf{X}^{\top})^{-1}\mathbf{X}$ as the projection matrix onto the row space of $\mathbf{X}$, denoted $\mathrm{R}(\mathbf{X})$. Then the vector of rewards for all arms, $\mathbf{Y}_{t}=(y_{1,t},\ldots,y_{K,t})$, is now decomposed as:
$\mathbf{Y}_{t}$ $=(\mathbf{X}^{\top}\boldsymbol{\theta}_{\star}^{(o)}+\mathbf{U}^{\top}\boldsymbol{\theta}_{\star}^{(u)})+\epsilon_{t}\boldsymbol{1}_{K}$
$=\mathbf{P}_{\mathbf{X}}(\mathbf{X}^{\top}\boldsymbol{\theta}_{\star}^{(o)}+\mathbf{U}^{\top}\boldsymbol{\theta}_{\star}^{(u)})+(\mathbf{I}_{K}-\mathbf{P}_{\mathbf{X}})(\mathbf{X}^{\top}\boldsymbol{\theta}_{\star}^{(o)}+\mathbf{U}^{\top}\boldsymbol{\theta}_{\star}^{(u)})+\epsilon_{t}\boldsymbol{1}_{K}$
$=\mathbf{X}^{\top}(\boldsymbol{\theta}_{\star}^{(o)}+(\mathbf{X}\mathbf{X}^{\top})^{-1}\mathbf{X}\mathbf{U}^{\top}\boldsymbol{\theta}_{\star}^{(u)})+(\mathbf{I}_{K}-\mathbf{P}_{\mathbf{X}})\mathbf{U}^{\top}\boldsymbol{\theta}_{\star}^{(u)}+\epsilon_{t}\boldsymbol{1}_{K},$ (4)
where the first and the second term are the projected rewards onto $\mathrm{R}(\mathbf{X})$ and $\mathrm{R}(\mathbf{X})^{\perp}$, the subspace of $\mathbb{R}^{K}$ perpendicular to $\mathrm{R}(\mathbf{X})$. We write the projected parameter as ${\boldsymbol{\mu}_{\star}^{(o)}}:=\boldsymbol{\theta}_{\star}^{(o)}+(\mathbf{X}\mathbf{X}^{\top})^{-1}\mathbf{X}\mathbf{U}^{\top}\boldsymbol{\theta}_{\star}^{(u)}$.
To handle the second term in Eq. 4, we consider a set of basis $\{\mathbf{b}_{1},\ldots,\mathbf{b}_{K-d}\}\in\mathrm{R}(\mathbf{X})^{\perp}$. Given the set, there exist coefficients ${\mu_{\star,1}^{(u)}},\dots,{\mu_{\star,K-d}^{(u)}}\in\mathbb{R}$ that express the reward projection as:
$(\mathbf{I}_{K}-\mathbf{P}_{\mathbf{X}})\mathbf{U}^{\top}\boldsymbol{\theta}_{\star}^{(u)}=\sum_{i=1}^{K-d}\mu_{\star,i}^{(u)}\mathbf{b}_{i}.$ (5)
We denote the number of nonzero coefficients:
$d_{h}(\mathbf{b}_{1},\ldots,\mathbf{b}_{K-d}):=|\{i\in[K-d]:\mu_{\star,i}^{(u)}\neq 0\}|.$ (6)
Note that $d_{h}=0$ for any basis $\{\mathbf{b}_{1},\ldots,\mathbf{b}_{K-d}\}$ when the latent feature space is included in the observed feature space, i.e., $\mathrm{R}(\mathbf{U})\subseteq\mathrm{R}(\mathbf{X})$, since it implies $(\mathbf{I}_{K}-\mathbf{P}_{X})\mathbf{U}^{\top}=\mathbf{0}_{K\times d_{u}}$. If $\mathrm{R}(\mathbf{U})\supseteq\mathrm{R}(\mathbf{X})$, then $d_{h}=K-d$ for any basis $\{\mathbf{b}_{1},\ldots,\mathbf{b}_{K-d}\}$. In other cases, the quantity $d_{h}$ depends on the choice of the basis $\{\mathbf{b}_{1},\ldots,\mathbf{b}_{K-d}\}$. Intuitively, $d_{h}$ tends to be small when $\mathrm{R}(\mathbf{X})$ and $\mathrm{R}(\mathbf{U})$ have a large intersection, while some choice of the basis may lead to a higher $d_{h}$ and consequently to a higher regret bound. Nevertheless, for any choice of the basis, our algorithm achieves $\widetilde{O}(\sqrt{(d+d_{h})T})$ regret without prior knowledge of $d_{h}$, which is smaller than $\widetilde{O}(\sqrt{KT})$ regret bound achieved by MAB algorithms ignoring features.
If we define $\boldsymbol{\mu}_{\star}$ as $[({\boldsymbol{\mu}_{\star}^{(o)}})^{\top},({\boldsymbol{\mu}_{\star}^{(u)}})^{\top}]^{\top}\in\mathbb{R}^{K}$, where $\boldsymbol{\mu}_{\star}^{(u)}=[{\mu_{\star,1}^{(u)}},\dots,{\mu_{\star,K-d}^{(u)}}]^{\top}$, then Eq. 4 becomes $\mathbf{Y}_{t}=[\mathbf{X}^{\top}\;\mathbf{b}_{1}\cdots\mathbf{b}_{K-d}]\boldsymbol{\mu}_{\star}+\epsilon_{t}\boldsymbol{1}_{K}$, implying that the reward for each $a\in[K]$ is:
$\begin{split}y_{a,t}&=\mathbf{e}_{a}^{\top}\mathbf{Y}\\ &=\mathbf{e}_{a}^{\top}[\mathbf{X}^{\top}\;\mathbf{b}_{1}\cdots\mathbf{b}_{K-d}]\boldsymbol{\mu}_{\star}+\epsilon_{t}\\ &=[{\mathbf{x}_{a}}\;\mathbf{e}_{a}^{\top}\mathbf{b}_{1}\cdots\mathbf{e}_{a}^{\top}\mathbf{b}_{K-d}]\boldsymbol{\mu}_{\star}+\epsilon_{t},\end{split}$ (7)
where $\mathbf{e}_{a}\in\mathbb{R}^{K}$ is a standard basis, with elements all zero except for $1$ in the $a$-th coordinate. With this modification, the rewards are now represented as a linear function of the augmented feature vectors, $\widetilde{\mathbf{x}}_{a}:=[\mathbf{x}_{a}^{\top}\;\mathbf{e}_{a}^{\top}\mathbf{b}_{1}\cdots\mathbf{e}_{a}^{\top}\mathbf{b}_{K-d}]^{\top}\in\mathbb{R}^{K}$, without any misspecification error. A toy example illustrating our strategy is shown in Figure 1.
Figure 1: Illustration comparing conventional linear bandit algorithms (left) and our approach (right) in estimating rewards of $K=3$ arms. Conventional algorithms use only observed features and find estimates within $\mathrm{R}(\mathbf{X})$ thus accumulating errors from unobserved features. However, our strategy projects the latent part of the reward onto the orthogonal complement of $\mathrm{R}(\mathbf{X})$, denoted by $\mathbf{b}_{1}^{\top}\widehat{\boldsymbol{\theta}}_{t}^{(u)}$, which enables reward estimation in $\mathbb{R}^{K}$. Note that $\hat{\boldsymbol{\theta}}_{t}$ is the estimator of the parameter for observed features.
$\mathbf{q}_{1}$$\mathbf{q}_{3}$$\mathbb{E}_{t-1}[\mathbf{Y}]$$\mathrm{R}(\mathbf{X})$$\mathbf{q}_{2}$$\mathbf{X}^{\top}\hat{\boldsymbol{\theta}}_{t}$
$\mathbf{q}_{1}$$\mathbf{q}_{3}$$\mathbb{E}_{t-1}[\mathbf{Y}]$$\mathbf{q}_{2}$($\mathrm{R}(\mathbf{X})$$\mathbf{b}_{1}^{\top}\widehat{\boldsymbol{\theta}}_{t}^{(u)}$$\mathbf{X}^{\top}\hat{\boldsymbol{\theta}}_{t}$
The dimension of the augmented feature vectors $\{\widetilde{\mathbf{x}}_{a}:a\in[K]\}$ is $K\geq d$. While applying SupLinUCB (Chu et al., 2011) yields $\widetilde{O}(\sqrt{KT})$ regret bound, it lacks practicality as it computes $\log T$ distinct batches and estimators requiring the knowledge of $T$ and, more critically, discards a significant portion of samples in each parameter update. We present an efficient algorithm that employs the doubly robust ridge estimator and achieves $\widetilde{O}(\sqrt{KT})$ regret bound (see Appendix A). However, when $K>d$ and $d_{u}=0$, the regret is high compared to the linear bandits with conventional features. Therefore, we propose a novel estimation strategy to avoid dependency on $K$ in the following section.
4.2 Doubly Robust Lasso Estimator
In Eq. 7, the parameter $\boldsymbol{\mu}_{\star}$ is sparse depending on the dimension of the latent features and the augmented basis vectors. Recall that $\boldsymbol{\mu}_{\star}^{(u)}$ are the coefficients to express the projection of the reward as represented in Eq. 5 and only $d_{h}$ basis vectors are required to express the projection of the reward; there are at most $d_{h}$ nonzero entries in $\boldsymbol{\mu}_{\star}^{(u)}$.
Let $\check{\boldsymbol{\mu}}^{L}_{t}$ denote the Lasso estimator for $\boldsymbol{\mu}_{\star}$ using augmented feature vectors:
$\begin{split}\check{\boldsymbol{\mu}}^{L}_{t}&:=\mathop{\mathrm{argmin}}_{\boldsymbol{\mu}}\sum_{\tau=1}^{t}(y_{a_{\tau}}-\widetilde{\boldsymbol{x}}_{a_{\tau}}^{\top}\boldsymbol{\mu})^{2}+2\sigma\sqrt{\frac{2t}{p}\log\frac{2Kt^{2}}{\delta}}\Big{\|}\big{(}\sum_{a\in[K]}\widetilde{\mathbf{x}}_{a}\widetilde{\mathbf{x}}_{a}^{\top}\big{)}^{\frac{1}{2}}\boldsymbol{\mu}\Big{\|}_{1}.\end{split}$ (8)
For the estimator in Eq. 8 to correctly identify the zero entries in $\boldsymbol{\mu}_{\star}$, the compatibility condition must hold (van de Geer and Bühlmann, 2009), which requires the minimum eigenvalue of the Gram matrix, $\lambda_{\min}\left(t^{-1}\sum_{s=1}^{t}\widetilde{\mathbf{x}}_{a_{s}}\widetilde{\mathbf{x}}_{a_{s}}^{\top}\right)$, to be positive. However, ensuring a sufficiently large minimum eigenvalue typically demands collecting a large number of exploration samples, thus increasing regret. Achieving this with fewer exploration samples remains a key challenge in bandit literature, as the minimum eigenvalue influences convergence rate of the estimator and, consequently, the regret bound (Kim et al., 2021; Soare et al., 2014).
We introduce a doubly robust (DR) estimator that employs the full feature Gram matrix $\sum_{s=1}^{t}\sum_{a=1}^{K}\widetilde{\mathbf{x}}_{a}\widetilde{\mathbf{x}}_{a}^{\top}$ instead of $\sum_{s=1}^{t}\widetilde{\mathbf{x}}_{a_{s}}\widetilde{\mathbf{x}}_{a_{s}}^{\top}$. The DR estimation originates from the statistical literature on missing data, where “doubly robust” means that the estimator is robust against errors in the estimation of both the observation probability and the response model. In bandits, at each decision round $t\in[T]$, only the reward of the selected arm is observed, while the $K-1$ unselected rewards are missing. Thus DR estimation is applied to impute these $K-1$ missing rewards and include corresponding $K-1$ feature vectors in the estimation. Since the observation probability is given by the policy (which is known to the learner), the DR estimator is robust against errors in the estimated rewards. While Kim and Paik (2019) proposed a DR Lasso estimator on IID features satisfying the compatibility condition, we propose another DR Lasso estimator that does not require the assumptions on features.
We improve the DR estimation by incorporating resampling and coupling methods. For each $t$, let $\mathcal{E}_{t}\subseteq[t]$ denote an exploration phase such that for $\tau\in\mathcal{E}_{t}$ the action $a_{\tau}$ is sampled uniformly over $[K]$. The $\mathcal{E}_{t}$ is constructed as $\mathcal{E}_{0}=\emptyset$ and $\mathcal{E}_{t}=\mathcal{E}_{t-1}\cup{t}$ if $|\mathcal{E}_{t}|\leq C_{e}\log(2Kt^{2}/\delta)$, otherwise $\mathcal{E}_{t}=\mathcal{E}_{t-1}$, where $C_{e}:=8(\sqrt{K}+p^{-1})^{2}p^{2}(1-p)^{-2}K^{2}\log\frac{2Kt^{2}}{\delta}$. In round $t$, when $t\notin\mathcal{E}_{t}$, the algorithm selects an action $a_{t}$ according to an $\epsilon_{t}$-greedy policy. Then, we generate a pseudo-action $\widetilde{a}_{t}$ from a multinomial distribution:
$\begin{split}&phi_{a_{t},t}:=\mathbb{P}(\widetilde{a}_{t}=a_{t}|a_{t})=p\quad\text{and}\quad\phi_{k,t}:=\mathbb{P}(\widetilde{a}_{t}=k|a_{t})=\frac{1-p}{K-1}\quad\forall k\in[K]\setminus\{a_{t}\},\end{split}$ (9)
where $p\in(1/2,1)$ is coupling probability set by the algorithm. To couple the policy of the actual action $a_{t}$ and the pseudo-action $\widetilde{a}_{t}$, we resample both of them until they match. This coupling yields a lower bound for the observation probability which reduces the variance of the DR pseudo-rewards in Eq. 10. Let $\mathcal{M}_{t}$ denote the event where $\widetilde{a}_{t}=a_{t}$ within a specified number of resamples. For given $\delta^{\prime}\in(0,1)$, we set the number of resamples as $\rho_{t}:=\log((t+1)^{2}/\delta^{\prime})/\log(1/p)$ so that event $\mathcal{M}_{t}$ occurs with probability at least $1-\delta^{\prime}/(t+1)^{2}$. Resampling allows the algorithm to explore further to find an action that balances between regret minimization and reward estimation.
This coupling replaces $\epsilon_{t}$ greedy policy with a multinomial distribution $\phi_{1,t},\ldots,\phi_{K,t}$. When we use DR estimation with $\epsilon_{t}$ greedy policy, the inverse probability $\epsilon_{t}^{-1}:=\sqrt{t}$ appears in the pseudo-reward (10), and thus the variance of the pseudo-reward explodes. Therefore, we couple the $\epsilon_{t}$ greedy policy with the multinomial distribution (9) to bound the inverse probability weight $\phi_{a,t}^{-1}=O(K)$.
With the pseudo-actions (coupled with the actual actions), we construct the unbiased pseudo-rewards for all $a\in[K]$,
$\widetilde{y}_{a,t}:={\widetilde{\mathbf{x}}_{a}}^{\top}\check{\boldsymbol{\mu}}_{t}^{L}+\frac{\mathbb{I}(\widetilde{a}_{t}=a)}{\phi_{a,t}}\big{(}y_{a,t}-\widetilde{\mathbf{x}}_{a}^{\top}\check{\boldsymbol{\mu}}_{t}^{L}\big{)},$ (10)
and note that $\check{\boldsymbol{\mu}}_{t}^{L}$ is the imputation estimator that fills in the missing rewards of unselected arms in round $t$.
For $a\neq\widetilde{a}_{t}$, i.e., an arm $a$ that is not selected in the round $t$, we impute the missing rewards using $\widetilde{\mathbf{x}}_{k}^{\top}\check{\boldsymbol{\mu}}_{t}^{L}$. For $a=\widetilde{a}_{t}$, however, the term $\mathbb{I}(\widetilde{a}_{t}=a)y_{a,t}/\phi_{a,t}$ calibrates the predicted reward to ensure the unbiasedness of the pseudo-rewards for all arms. Given that $\mathbb{E}_{\widetilde{a}_{t}}[\mathbb{I}(\widetilde{a}_{t}=a)]=\mathbb{P}(\widetilde{a}_{t}=a)=\phi_{a,t}$, taking the expectation over $\widetilde{a}_{t}$ on both sides of Eq. 10 gives $\mathbb{E}_{\widetilde{a}_{t}}[\widetilde{Y}_{a,t}]=\mathbb{E}_{t-1}[y_{a,t}]=\widetilde{\mathbf{x}}_{a}^{\top}{{\boldsymbol{\mu}}_{\star}}$ for all $a\in[K]$. Although the estimate $\widetilde{\mathbf{x}}_{a}^{\top}\check{\boldsymbol{\mu}}_{t}$ may have high error, it is multiplied by the mean-zero random variable $(1-\mathbb{I}(\widetilde{a}_{t}=a)/\phi_{a,t})$, making the pseudo-rewards robust to the error in $\widetilde{\mathbf{x}}_{a}^{\top}\check{\boldsymbol{\mu}}_{t}$. The pseudo-rewards can only be computed when $\widetilde{a}_{t}=a_{t}$, so they are used in rounds when the chosen action $a_{t}$ and the pseudo-action $\widetilde{a}_{t}$ match, indicated by the event $\mathcal{M}_{t}$. Since $\mathcal{M}_{t}$ occurs with high probability, we can compute the pseudo-rewards for almost all rounds.
Our DR Lasso estimator, which incorporates pseudo-rewards for estimation, is defined as follows:
$\begin{split}\widehat{\boldsymbol{\mu}}^{L}_{t}&:=\mathop{\mathrm{argmin}}_{\boldsymbol{\mu}}\sum_{\tau=1}^{t}\mathbb{I}(\mathcal{M}_{\tau})\sum_{a\in[K]}\left(\widetilde{y}_{a,\tau}-\widetilde{\mathbf{x}}_{a}^{\top}{\boldsymbol{\mu}}\right)^{2}+\frac{10\sigma}{p}\sigma\sqrt{2t\log\frac{2Kt^{2}}{\delta}}\Big{\|}\big{(}\sum_{a\in[K]}\widetilde{\mathbf{x}}_{a}\widetilde{\mathbf{x}}_{a}^{\top}\big{)}^{1/2}\boldsymbol{\mu}\Big{\|}_{1},\end{split}$ (11)
and the following theorem provides a theoretical guarantee that this estimator converges across all arms after several exploration rounds.
Theorem 2 (Consistency of the DR Lasso estimator).
Let $d_{h}$ denote the dimension of the projected latent rewards defined in Eq. 6. Then for all round $t$ such that $t\geq|\mathcal{E}_{t}|$, with probability at least $1-2\delta/t^{2}$,
$\max_{a\in[K]}|\widetilde{\mathbf{x}}_{a}(\widehat{\boldsymbol{\mu}}_{t}-\boldsymbol{\mu}_{\star})|\leq\frac{20\sigma}{p}\sqrt{\frac{2(d+d_{h})\log\frac{2Kt^{2}}{\delta}}{t}}.$ (12)
Although the DR Lasso estimator leverages $K$-dimensional feature vectors, its error bound depends only logarithmically on $K$. Usually, such fast convergence is made possible under classical regularity conditions, such as the compatibility condition and the restrictive minimum eigenvalue condition (Bühlmann and van de Geer, 2011; van de Geer and Bühlmann, 2009). Also, the existing Lasso-based bandit approaches (Bastani and Bayati, 2020; Kim and Paik, 2019; Oh et al., 2021; Ariu et al., 2022; Chakraborty et al., 2023; Lee et al., 2025) impose these conditions directly on the feature vectors. In contrast, our approach does not require this assumption, since our augmented features are orthogonal vectors in $\mathrm{R}(\mathbf{X})^{\perp}$ and their average Gram matrix satisfies $\lambda_{\min}(\sum_{a\in[K]}\widetilde{\mathbf{x}}_{a}\widetilde{\mathbf{x}}_{a}^{\top})\geq\min\{1,\lambda_{\min}(\sum_{a\in[K]}\mathbf{x}_{a}\mathbf{x}_{a}^{\top})\}$. Thus, the convergence rate has only $\sqrt{\log K}$ rate in terms of $K$.
The consistency is proved by bounding the two components of the error in the pseudo-rewards defined in (10): (i) the noise of the reward and (ii) the error of the imputation estimator $\check{\mu}_{t}$. Since (i) is sub-Gaussian, it can be bounded using martingale inequalities. For (ii), the imputation error $\widetilde{x}_{a}^{\top}(\check{\mu}^{L}_{t}-\mu_{\star})$ is multiplied by the mean-zero random variable $\left(1-\frac{\mathbb{I}(\widetilde{a}_{t}=a)}{\phi_{a,t}}\right)$ and thus it can be bounded by $\|\check{\mu}^{L}_{t}-\mu_{\star}\|_{1}/\sqrt{t}$.
Algorithm 1 Robust to Latent Feature (RoLF)
1:INPUT: features $\{\mathbf{x}_{a}:a\in[K]\}$, coupling probability $p\in(1/2,1)$, confidence parameter $\delta>0$.
2:Initialize $\widehat{\boldsymbol{\mu}}_{0}=\mathbf{0}_{K}$, the exploration phase $\mathcal{E}_{t}=\emptyset$ and the exploration factor $C_{\mathrm{e}}:=8(\sqrt{K}+p^{-1})^{2}p^{2}(1-p)^{-2}K^{2}\log\frac{2Kt^{2}}{\delta}$.
3:Find orthogonal basis $\mathbf{b}_{1},\ldots,\mathbf{b}_{K-d}$ in $\mathrm{R}(\mathbf{X})^{\perp}$ to construct $\{{\widetilde{\mathbf{x}}_{a}}:a\in[K]\}$
4:for $t=1,\ldots,T$ do
5:     if $|\mathcal{E}_{t}|\leq C_{\mathrm{e}}\log(2Kt^{2}/\delta)$ then
6:         Randomly sample $a_{t}$ uniformly over $[K]$ and $\mathcal{E}_{t}=\mathcal{E}_{t-1}\cup\{t\}$.
7:     else
8:         Compute $\widehat{a}_{t}:=\arg\max_{a\in[K]}{\widetilde{\mathbf{x}}_{a}}^{\top}\widehat{\boldsymbol{\mu}}^{L}_{t-1}$
9:         while $\widetilde{a}_{t}\neq a_{t}$ and $\text{count}\leq\rho_{t}$ do
10:              Sample $a_{t}$ with $\mathbb{P}(a_{t}=\widehat{a}_{t})=1-(t^{-1/2})$ and $\mathbb{P}(a_{t}=k)=t^{-1/2}/(K-1),\;\forall k\neq\widehat{a}_{t}$.
11:              Sample $\widetilde{a}_{t}$ according to  Eq. 9.
12:              $\text{count}=\text{count}+1$
13:     Play $a_{t}$ and observe $y_{a_{t},t}$.
14:     if $\widetilde{a}_{t}\neq a_{t}$ then
15:         Set $\widehat{\boldsymbol{\mu}}^{L}_{t}:=\widehat{\boldsymbol{\mu}}^{L}_{t-1}$
16:     else
17:         Update $\widehat{\boldsymbol{\mu}}^{L}_{t}$ following Eq. 11 with $\widetilde{y}_{a,t}$ and update $\check{\boldsymbol{\mu}}^{L}_{t}$ following Eq. 8.
5 Proposed Algorithm and Regret Analysis
In this section, we propose Algorithm 1, a novel estimation-based algorithm for handling partially observable features.
5.1 Robust to Latent Features (RoLF) Algorithm
In the initialization step, when the observed features are given, our algorithm finds a set of orthogonal basis $\{\mathbf{b}_{1},\dots,\mathbf{b}_{K-d}\}\in\mathrm{R}(\mathbf{X})^{\perp}$ to augment each observed features. After the forced-exploration phase, the algorithm computes the candidate action, denoted by $\widehat{a}_{t}$, and then resample both $\widetilde{a}_{t}$ and $a_{t}$ until they match. When the resampling phase ends, and the agent selects $a_{t}$ and observes $y_{a_{t},t}$. If they match within $\rho_{t}:=\log((t+1)^{2}/\delta^{\prime})/\log(1/p)$, the algorithm updates both the imputation and main estimators; otherwise neither estimator is updated.
The proposed algorithm does not require the knowledge of the dimension of the latent features $d_{u}$ and the dimension of the projected rewards from latent feature space onto the $\mathrm{R}(\mathbf{X})^{\perp}$. Although we present the algorithm on fixed feature vectors, the algorithm applies to arbitrary feature vectors that changes over time by updating the orthogonal basis.
5.2 Regret Analysis
We analyze the regret bound of RoLF using Lasso estimators, as stated in the following theorem:
Theorem 3 (Regret bound for Lasso RoLF).
Let $d_{h}$ denote the dimension of the projected latent rewards defined in Eq. 6. Then for $\delta\in(0,1)$ and $p\in(1/2,1)$, the expected cumulative regret of the proposed algorithm is bounded by
$\begin{split}\mathbb{E}[\mathrm{Reg}(T)]&\leq 10\delta\log T+\frac{16K^{2}(\sqrt{K}+p^{-1})^{2}}{(1-p)^{2}}+\frac{4p\sqrt{T}}{K-1}\frac{\log\frac{(T+1)^{2}}{\delta}}{\log(1/p)}+\frac{80\sigma}{p}\sqrt{2(d+d_{h})T\log\frac{2KT^{2}}{\delta}},\end{split}$ (13)
To the best of our knowledge, Theorem 3 is the first regret bound sublinear in $T$ for the latent features without any structural assumption. Note that the number of rounds for the exploration phase is $O(K^{3}\log KT)$, which is only logarithmic in the horizon $T$. The factor $K^{3}$ is not reducible since the algorithm must estimate all $K$ biases from the missing features. Using the Gram matrix with full feature vectors, $\sum_{a=1}^{K}\widetilde{\mathbf{x}}_{a}\widetilde{\mathbf{x}}_{a}^{\top}$ in combination with DR estimation reduces the exploration phase time from $O(K^{4}\log KT)$ to $O(K^{3}\log KT)$, reducing the complexity by a factor of $K$. The convergence rate in the last term is proportional to $\sqrt{d+d_{h}}$ rather than $\sqrt{K}$, as shown in Eq. 12. Thus, our regret bound is $O(\sqrt{(d+d_{h})T\log KT})$.
6 Numerical Experiments
In this experiment, we simulate and compare two versions of our algorithm, presented in Algorithm 1 and Algorithm 2 (Appendix A), with linear bandit algorithms that use only observed features: LinUCB (Li et al., 2010; Chu et al., 2011) and LinTS. These algorithms use the UCB and Thompson sampling methods, respectively, when the reward is modeled as a linear function of the features. Additionally, since our algorithm incorporates DR estimation with the Lasso estimator, we include DRLasso (Kim and Paik, 2019) in the comparison as well. To further evaluate the performance of our algorithm in scenarios where latent features are expected but ignored, we also compare it with UCB($\delta$) (Lattimore and Szepesvári, 2020), an MAB algorithm without features.
For the simulation environment, we generate true features $\mathbf{z}_{a}$ for each arm $a\in[K]$ from $\mathcal{N}(\mathbf{0},\mathbf{I}_{d_{z}})$ and subsample $d$ elements to construct $\mathbf{x}_{a}$. Orthogonal basis vectors $\{\mathbf{b}_{1},\dots,\mathbf{b}_{K-d}\}$ are derived via singular value decomposition (SVD) on the observed feature matrix $\mathbf{X}$, ensuring orthogonality to $\mathrm{R}(\mathbf{X})$. We augment $\mathbf{X}$ with the basis vectors via linear concatenation. Rewards are generated by sampling the unknown parameter $\boldsymbol{\theta}_{\star}\in\mathbb{R}^{k}$ from $\mathrm{Unif}(-1/2,1/2)$. The hyperparameter $p$, for the sampling distribution of $\widetilde{a}_{t}$, is set to $0.6$ (see Eq. 9). The confidence parameter $\delta$ is $10^{-4}$, and the total decision horizon is $T=1200$. To address both partial and full observability, $d_{z}\geq d$ is used, and we run 5 independent experiments. We compare the algorithms across three scenarios:
(a) $d=1$
(b) $d=\left\lfloor\frac{d_{z}}{2}\right\rfloor$
(c) $d=d_{z}-1$
Figure 2: Cumulative regrets of the algorithms in comparison for scenario 1 ($K=50,d_{z}=31$).
(a) $K=d=20$
(b) $K=d=30$
(c) $K=d=40$
Figure 3: Cumulative regrets of the algorithms in comparison for scenario 2 ($d_{z}=60$).
(a) $K=15$, $d=30$
(b) $K=20$, $d=40$
(c) $K=30$, $d=60$
Figure 4: Cumulative regrets of the algorithms in comparison for scenario 3 ($d_{z}=d$).
Scenario (i).
We examine algorithm performance as $d$, the number of observed elements, varies to assess the impact of observability. With $K=50$ arms and $d_{z}=31$, we compare results for $d=1$, $\lfloor d_{z}/2\rfloor=15$, and $d_{z}-1=30$. Figure 2 presents the results, showing that our algorithm consistently outperforms others in regret and robustness. In contrast, LinUCB, LinTS, and DRLasso show significant dependence on the number of observed features, with performance deteriorating and variability increasing as observability decreases.
Scenario (ii).
In this scenario, the number of arms is equal to the dimension of the observed features, i.e., $K=d$. The main objective of this experiment is to demonstrate that our algorithm remains robust to changes in the number of arms $K$, unlike MAB algorithms that disregard the observed features. Specifically, we compare the algorithms with $K$ set to 20, 30, and 40, and $d_{z}=60$ for all cases. Figure 3 shows the results for this scenario. As we can observe, the performance of UCB($\delta$) worsens as the number of arms increases in each environment. In contrast, our algorithm shows better performance in terms of both the level of regret and robustness.
Scenario (iii).
We evaluate performance when the number of arms is less than the dimension of observed features, setting $d=2K$ and varying $K$ as 15, 20, and 30, with $d_{z}=d$. Before using the features in our algorithms, we apply singular value decomposition (SVD) for dimensionality reduction. Figure 4 shows that our algorithm performs well even in extreme cases. By applying dimension reduction through SVD, our algorithm remains applicable even when the matrix of feature vectors is not full rank. Furthermore, the results suggest that our algorithm demonstrates superior performance even in the absence of partial observability.
7 Conclusion
In this paper, we addressed the challenges due to partially observable features in linear bandit problems. We showed that algorithms relying solely on observed features may suffer linear regret due to model misspecification and introduced RoLF, a novel algorithm that incorporates latent features without prior knowledge of them. Our algorithm improves the regret bound over traditional methods that ignore latent features, with numerical experiments supporting the theoretical results. Moving forward, several open directions remain for future work. First, while the linear bandit framework allows our feature augmentation strategy in Section 4.1 to be interpreted as another linear bandit problem, extending this approach to more general reward functions, such as generalized linear models, would be interesting. Furthermore, as noted in Section 3.2, we assumed the latent portion of reward to be linear with respect to the unobserved features. However, the unobserved reward can be viewed as an exogenous factor affecting the agent’s learning. This perspective enables extending the unobserved reward to a general function class without structural assumptions.
Appendix A Robust to Latent Feature Algorithm with Ridge Estimator
Algorithm 2 Robust to Latent Feature with Ridge Estimator (RoLF-Ridge)
1:INPUT: features $\{\mathbf{x}_{a}:a\in[K]\}$, coupling probability $p\in(1/2,1)$, confidence parameter $\delta>0$.
2:Initialize $\widehat{\boldsymbol{\mu}}_{0}=\mathbf{0}_{K}$, the exploration phase $\mathcal{E}_{t}=\emptyset$ and the exploration factor $C_{\mathrm{e}}:=32(1-p)^{-2}K^{2}$.
3:Find orthogonal basis $\mathbf{b}_{1},\ldots,\mathbf{b}_{K-d}$ in $\mathrm{R}(\mathbf{X})^{\perp}$ to construct $\{{\widetilde{\mathbf{x}}_{a}}:a\in[K]\}$
4:for $t=1,\ldots,T$ do
5:     if $|\mathcal{E}_{t}|\leq C_{\mathrm{e}}\log(2Kt^{2}/\delta)$ then
6:         Randomly sample $a_{t}$ uniformly over $[K]$ and $\mathcal{E}_{t}=\mathcal{E}_{t-1}\cup\{t\}$.
7:     else
8:         Compute $\widehat{a}_{t}:=\arg\max_{a\in[K]}{\widetilde{\mathbf{x}}_{a}}^{\top}\widehat{\boldsymbol{\mu}}^{R}_{t-1}$
9:         while $\widetilde{a}_{t}\neq a_{t}$ and $\text{count}\leq\rho_{t}$ do
10:              Sample $a_{t}$ with $\mathbb{P}(a_{t}=\widehat{a}_{t})=1-(t^{-1/2})$ and $\mathbb{P}(a_{t}=k)=t^{-1/2}/(K-1),\;\forall k\neq\widehat{a}_{t}$.
11:              Sample $\widetilde{a}_{t}$ according to  Eq. 9.
12:              $\text{count}=\text{count}+1$
13:     Play $a_{t}$ and observe $y_{a_{t},t}$.
14:     if $\widetilde{a}_{t}\neq a_{t}$ then
15:         Set $\widehat{\boldsymbol{\mu}}^{R}_{t}:=\widehat{\boldsymbol{\mu}}^{R}_{t-1}$
16:     else
17:         Update $\widehat{\boldsymbol{\mu}}_{t}^{R}$ following Eq. 14 with $\widetilde{y}_{a,t}$ and update $\check{\boldsymbol{\mu}}_{t}^{R}$ following Eq. 15.
Our Doubly robust (DR) ridge estimator is defined as follows:
$\widehat{\boldsymbol{\mu}}^{R}_{t}:=\left(\sum_{\tau=1}^{t}\mathbb{I}(\mathcal{M}_{\tau})\sum_{a\in[K]}{\widetilde{\mathbf{x}}_{a}}{\widetilde{\mathbf{x}}_{a}}^{\top}+I_{K}\right)^{-1}\left(\sum_{\tau=1}^{t}\mathbb{I}(\mathcal{M}_{\tau})\sum_{a\in[K]}{\widetilde{\mathbf{x}}_{a}}\widetilde{y}_{a,\tau}\right),$ (14)
where $\widetilde{y}_{a,\tau}$ is the DR pseudo reward:
$\widetilde{y}_{a,t}:={{\widetilde{\mathbf{x}}_{a}}}^{\top}\check{\boldsymbol{\mu}}_{t}^{R}+\frac{\mathbb{I}(\widetilde{a}_{t}=a)}{\phi_{a,t}}\big{(}y_{a,t}-{\widetilde{\mathbf{x}}_{a}}^{\top}\check{\boldsymbol{\mu}}_{t}^{R}\big{)},$
and the imputation estimator $\check{\boldsymbol{\mu}}^{R}_{t}$ is defined as
$\check{\boldsymbol{\mu}}^{R}_{t}:=\left(\sum_{\tau=1}^{t}{\widetilde{\mathbf{x}}_{a_{\tau}}}{\widetilde{\mathbf{x}}_{a_{\tau}}}^{\top}+p\mathbf{I}_{K}\right)^{-1}\left(\sum_{\tau=1}^{t}{\widetilde{\mathbf{x}}_{a_{\tau}}}y_{a_{\tau},\tau}\right).$ (15)
The following theorem shows that this Ridge estimator is consistent, meaning it converges to the true parameter ${\boldsymbol{\mu}}_{\star}$ with high probability as the agent interacts with the environment.
Theorem 4 (Consistency of the DR Ridge estimator).
For each $t$, let $\mathcal{E}_{t}\subseteq[t]$ denote an exploration phase such that for $\tau\in\mathcal{E}_{t}$ the action $a_{\tau}$ is sampled uniformly over $[K]$. Then for all round $t$ such that $|\mathcal{E}_{t}|\geq 32(1-p)^{-2}K^{2}\log(2Kt^{2}/\delta)$, with probability at least $1-3\delta$,
$\max_{a\in[K]}|{\widetilde{\mathbf{x}}_{a}}^{\top}(\widehat{\boldsymbol{\mu}}^{R}_{t}-{\boldsymbol{\mu}}_{\star})|\leq\frac{2}{\sqrt{t}}\left(\frac{\sigma}{p}\sqrt{K\log\frac{t+1}{\delta}}+\sqrt{K}\right).$
With $|\mathcal{E}_{t}|=O(K^{2}\log Kt)$ number of exploration, the DR Ridge estimator achieves $O(\sqrt{K/t})$ convergence rate over all $K$ rewards. This is possible because the DR pseudo-rewards defined in Eq. 10 impute the missing rewards for all arms $a\in[K]$ using ${\widetilde{\mathbf{x}}_{a}}^{\top}\check{\boldsymbol{\mu}}_{t}$, based on the samples collected during the exploration phase, $\mathcal{E}_{t}$. With this convergence guarantee, we establish a regret bound for RoLF-Ridge, which is the adaptation of Algorithm 1 using the Ridge estimator.
Theorem 5 (Regret bound for Ridge RoLF).
For $\delta\in(0,1)$, the expected cumulative regret of the proposed algorithm using DR Ridge estimator is bounded by
$\mathrm{Reg}(T)\leq 6\delta\log T+\frac{2p\sqrt{T}}{K-1}\frac{\log\frac{(T+1)^{2}}{\delta}}{\log(1/p)}+\frac{32K^{2}}{(1-p)^{2}}\log\frac{2dT^{2}}{\delta}+8\sqrt{KT}\left(\frac{\sigma}{p}\sqrt{\log\frac{T^{2}}{\delta}}+1\right).$
The first and second terms come from the distribution of $a_{t}$ which is a combination of the $1-t^{-1/2}$-greedy policy and resampling up to $\rho_{t}:=\log((t+1)^{2}/\delta)/\log(1/p)$ trials. The third term is determined by the size of the exploration set, $\mathcal{E}_{t}$, while the last term arises from the estimation error bounded by the DR estimator as described in Theorem 4. The hyperparameter $p\in(1/2,1)$ balances the size of the exploration set in the third term and the estimation error in the last term. Overall, the regret is $O(\sqrt{KT\log T})$, which shows a significant improvement compared to the regret lower bound in Theorem 1 for any linear bandit algorithms that do not account for unobserved features and unobserved rewards.
Appendix B A Modified Algorithm for Time-Varying Observed Features
In this section, we propose an algorithm for linear bandits with partially observable features under the setting where the observed features vary over time.
B.1 Problem Formulation
Let $\mathbf{x}_{1,t},\ldots,\mathbf{x}_{K,t}$ denote the observed features and $\mathbf{u}_{1},\ldots,\mathbf{u}_{K}$ denote the unobserved features. Now the observed features arbitrary changes over $t$ but the unobserved features are fixed over time. When the algorithms selects an arm $a_{t}$, the reward is
$y_{a_{t},t}=\langle\mathbf{x}_{a_{t},t},\boldsymbol{\theta}_{\star}^{(o)}\rangle+\langle\mathbf{u}_{a_{t}},\boldsymbol{\theta}_{\star}^{(u)}\rangle+\epsilon_{t},$
where the $\epsilon_{t}$ is the Sub-Gaussian noise that follows Assumption 1. The expected reward of each arm is sta over time, where MAB algorithms without using features are applicable to achieve $\widetilde{O}(\sqrt{KT})$ regret bound. When the observed features vary over time, the expected reward of each arm $\mathbb{E}[y_{a_{t}},t]=\langle\mathbf{x}_{a_{t},t},\boldsymbol{\theta}_{\star}^{(o)}\rangle+\langle\mathbf{u}_{a_{t}},\boldsymbol{\theta}_{\star}^{(u)}\rangle$ also arbitrarily changes over time and MAB algorithms suffer regret linear in $T$. To our knowledge, there is no other work that address this challenging setting.
B.2 Proposed Method: Orthogonal Basis Augmentation
We address the problem by augmenting Euclidean basis $\mathbf{e}_{1},\ldots,\mathbf{e}_{K}$ in $\mathbb{R}^{K}$ to estimate bias caused by the unobserved features. Let $\widetilde{\mathbf{x}}_{a,t}:=\mathbf{e}_{a}^{\top}[\mathbf{X}_{t}\;\mathbf{e}_{1}\;\cdots\;\mathbf{e}_{K}]\in\mathbb{R}^{d+K}$ and let $\Delta_{a}:=\langle\mathbf{u}_{a},\theta_{\star}^{(u)}\rangle$ denote the bias stems from the latent features. Then,
$y_{a,t}$ $=\langle\mathbf{x}_{a,t}^{\top}\boldsymbol{\theta}_{\star}^{(o)}\rangle+\langle\mathbf{u}_{a,t}^{\top}\boldsymbol{\theta}_{\star}^{(u)}\rangle+\epsilon_{a,t}$
$=\langle\mathbf{e}_{a}^{\top}[\mathbf{X}_{t}\mathbf{e}_{1}\cdots\mathbf{e}_{K}],[\boldsymbol{\theta}_{\star}^{(o)}\Delta_{1}\cdots\Delta_{K}]\rangle+\epsilon_{a,t}.$
Therefore, applying the RoLF-Ridge algorithm to the new features $\widetilde{\mathbf{x}}_{a,t}:=\mathbf{e}_{a}^{\top}[\mathbf{X}_{t}\mathbf{e}_{1}\cdots\mathbf{e}_{K}]$ yields the following regret bound.
Theorem 6 (Regret bound for Ridge-RoLF-V with time varying observed features).
If observed features are vary over time, for $\delta\in(0,1)$, the expected cumulative regret of the proposed algorithm Ridge-RoLF-V using DR Ridge estimator is bounded by
$\begin{split}\mathrm{Reg}(T)\leq&6\delta\log T+\frac{2p\sqrt{T}}{d+K-1}\frac{\log\frac{(T+1)^{2}}{\delta}}{\log(1/p)}+\frac{32(K+d)^{2}}{(1-p)^{2}}\log\frac{2(K+d)T^{2}}{\delta}\\ &+8\sqrt{(d+K)T}\left(\frac{\sigma}{p}\sqrt{\log\frac{T^{2}}{\delta}}+1\right).\end{split}$
The proof is similar to that in Theorem 5 and we omit the proof. The rate of the regret bound is $\widetilde{O}(\sqrt{(d+K)T})$ and, to our knowledge, this is the first sublinear regret bound for the partially observable linear bandits (as well as misspecified linear bandits) with arbitral time-varying observed features.
Appendix C Missing Proofs
C.1 Proof of Theorem 1
Throughout this paper, we consider a bandit problem where the agent observes only a subset of the reward-generating feature vector and cannot access or estimate the unobserved portion. If the agent uses online decision-making algorithms that rely solely on observed features, as defined in Definition 1, the resulting issue can be interpreted as a model misspecification. Therefore, in this theorem, we present a problem instance where “misspecified” algorithms, considering only observed features, may incur regret that grows linearly in $T$.
Following the statement of Theorem 1, we assume that $d=d_{u}=1$, which means $d_{z}=2$. Given the true feature set $\mathcal{Z}=\{[1,3]^{\top},[2,19/4]^{\top}\}$, let the first element of each vector is observed to the agent; while the second element remains unobserved. This results in $\mathbf{x}_{1}=x_{1}=1$, $\mathbf{x}_{2}=x_{2}=2$, $\mathbf{u}_{1}=u_{1}=3$ and $\mathbf{u}_{2}=u_{2}=19/4$. We set the true parameter as $\boldsymbol{\theta}_{\star}\in\mathbb{R}^{2}=[2,-1]^{\top}$, meaning $\boldsymbol{\theta}_{\star}^{(o)}=\theta_{\star}^{(o)}=2$ and $\boldsymbol{\theta}_{\star}^{(u)}=\theta_{\star}^{(u)}=-1$. Using the reward function from Section 3.2 and considering Assumption 1, the expected reward for each arm is given by
$\gamma_{i}:=\mathbb{E}[y_{i}]=\mathbf{z}_{i}^{\top}\theta_{\star}=x_{i}\theta_{\star}^{(o)}+u_{i}\theta_{\star}^{(u)}\quad\forall\ i\in\{1,2\}.$
Plugging the values in, the true mean reward for each arm is directly computed as $\gamma_{1}=2-3=-1$ and $\gamma_{2}=4-19/4=-3/4$, which satisfies the assumption that its absolute value does not exceed 1 (Section 3.2), and since $\gamma_{1}<\gamma_{2}$, the arm 2 is the optimal action.
For brevity, we denote the latent reward components as $g_{1}:=u_{1}\theta_{\star}^{(u)}$ and $g_{2}:=u_{2}\theta_{\star}^{(u)}$, yielding that $g_{1}=-3$ and $g_{2}=-19/4$. Note that since $\gamma_{2}\neq 2\gamma_{1}$ and $|{g_{i}}|\geq 3>0$ for all $i\in\{1,2\}$, thus our problem setup satisfies the “large deviation” criterion in Definition 1 and Theorem 2 of Ghosh et al. [2017], by letting $l=3$ and $\beta=0$. Consequently, by applying the theorem, it follows that using OFUL in this problem instance results in linear regret with respect to $T$, i.e., $\Omega(T)$. Inspired by this theorem, we demonstrate that LinTS encounters the same issue.
For each round $t\in[T]$, LinTS estimates the true parameter using the ridge estimator, given by:
$\begin{split}\hat{\boldsymbol{\theta}}_{t}&=(\mathbf{X}_{t}^{\top}\mathbf{X}_{t}+\lambda\mathbf{I}_{d})^{-1}(\mathbf{X}_{t}^{\top}\mathbf{Y}_{t})\\ &=(\mathbf{X}_{t}^{\top}\mathbf{X}_{t}+\lambda\mathbf{I}_{d})^{-1}(\mathbf{X}_{t}^{\top}(\mathbf{X}_{t}\boldsymbol{\theta}_{\star}^{(o)}+\mathbf{g}_{t}+\boldsymbol{\epsilon}_{t}))\\ &=\boldsymbol{\theta}_{\star}^{(o)}-\lambda\mathbf{V}_{t}^{-1}\boldsymbol{\theta}_{\star}^{(o)}+\mathbf{V}_{t}^{-1}\mathbf{X}_{t}^{\top}\mathbf{g}_{t}+\mathbf{V}_{t}^{-1}\mathbf{X}_{t}^{\top}\boldsymbol{\epsilon}_{t},\end{split}$ (16)
where $\mathbf{X}_{t}:=(\mathbf{x}_{a_{1}}^{\top},\dots,\mathbf{x}_{a_{t}}^{\top})\in\mathbb{R}^{t\times d}$ is a matrix containing features chosen up to round $t$, $\mathbf{Y}_{t}:=(y_{a_{1}},\dots,y_{a_{t}})\in\mathbb{R}^{t}$ is a vector of observed rewards, and $\boldsymbol{\epsilon}_{t}:=(\epsilon_{1},\dots,\epsilon_{t})\in\mathbb{R}^{t}$ contains noise attached to each reward. Unlike a typical ridge estimator, here the term $\mathbf{g}_{t}:=(g_{a_{1}},\dots,g_{a_{t}})\in\mathbb{R}^{t}$, the vector containing the latent portion of observed rewards is introduced due to model misspecification. Note that $\mathbf{V}_{t}:=(\mathbf{X}_{t}^{\top}\mathbf{X}_{t}+\lambda\mathbf{I}_{d})\succ 0$.
For this problem instance, since $d=1$Eq. 16 is equivalent to:
$\hat{\theta}_{t}=\theta_{\star}^{(o)}-\frac{\theta_{\star}^{(o)}}{\sum_{\tau=1}^{t}x_{a_{\tau}}^{2}+1}+\frac{\sum_{\tau=1}^{t}x_{a_{\tau}}g_{a_{\tau}}}{\sum_{\tau=1}^{t}x_{a_{\tau}}^{2}+1}+\frac{\sum_{\tau=1}^{t}x_{a_{\tau}}\epsilon_{\tau}}{\sum_{\tau=1}^{t}x_{a_{\tau}}^{2}+1},$
where we assume $\lambda=1$. Note also that we denote $\hat{\boldsymbol{\theta}}_{t}$ and $\boldsymbol{\theta}_{\star}^{(o)}$ by $\hat{\theta}_{t}$ and $\theta_{\star}^{(o)}$, respectively, since both are scalars. Hence, the estimation error is computed as:
$\hat{\theta}_{t}-\theta_{\star}^{(o)}=-\frac{\theta_{\star}^{(o)}}{\sum_{\tau=1}^{t}x_{a_{\tau}}^{2}+1}+\frac{\sum_{\tau=1}^{t}x_{a_{\tau}}g_{a_{\tau}}}{\sum_{\tau=1}^{t}x_{a_{\tau}}^{2}+1}+\frac{\sum_{\tau=1}^{t}x_{a_{\tau}}\epsilon_{\tau}}{\sum_{\tau=1}^{t}x_{a_{\tau}}^{2}+1}.$ (17)
Let $N_{1}$ and $N_{2}$ denote the number of times arms 1 and 2 have been played up to round $t$, respectively. This implies that $N_{1}+N_{2}=t$. Then, for the numerator of the second term, since
$\sum_{\tau=1}^{t}x_{a_{\tau}}g_{a_{\tau}}=(\underbrace{g_{1}+\cdots+g_{1}}_{N_{1}}+\underbrace{2g_{2}+\cdots+2g_{2}}_{N_{2}})=g_{1}N_{1}+2g_{2}N_{2},$
we can observe that
$\sum_{\tau=1}^{t}x_{a_{\tau}}g_{a_{\tau}}$ $=g_{1}N_{1}+2g_{2}N_{2}$
$\geq\bar{g}N_{1}+2\bar{g}N_{2}$
$=\bar{g}N_{1}+2\bar{g}(t-N_{1})$
$=2\bar{g}t-\bar{g}N_{1}$
$\geq\bar{g}t\qquad\qquad(\because N_{1}\leq t)$
$=-\frac{19}{4}t,$
where $\bar{g}=\min\{g_{1},g_{2}\}=-19/4$, which implies that $\sum_{\tau=1}^{t}x_{a_{\tau}}g_{a_{\tau}}=\Theta(t)$. For the denominator, $\sum_{\tau=1}^{t}x_{a_{\tau}}^{2}+1$, as it grows at a rate of $O(t)$, implying that the second term in Eq. 17 is $\Theta(1)$, and that $\hat{\theta}_{t}$ is not consistent since it does not converge to $\theta_{\star}^{(o)}$ as $t\to\infty$.
For arm 2, which is optimal, to be selected in round $t+1$ under LinTS, the condition $x_{2}\widetilde{\theta}_{t}\geq x_{1}\widetilde{\theta}_{t}$ must hold, where $\widetilde{\theta}_{t}\sim\mathcal{N}\left(\hat{\theta}_{t},\frac{v^{2}}{\sum_{\tau=1}^{t}x_{a_{\tau}}^{2}+1}\right)$. Given the assumptions that $x_{1}=1$ and $x_{2}=2$, arm 2 is selected whenever $\widetilde{\theta}_{t}\geq 0$. Thus, for arm 1 to be chosen, we require $\widetilde{\theta}_{t}<0$. We will show that the probability of $\widetilde{\theta}_{t}<0$ does not diminish sufficiently to be ignored even the agent plays sufficiently large amount of time. To clarify, let us define two events $E_{\widetilde{\theta}}:=\{\widetilde{\theta}_{t}\geq 0\}$ and $E_{\hat{\theta}}:=\{\hat{\theta}_{t}\geq 0\}$.
We revisit Eq. 16 as follows:
$\hat{\theta}_{t}$ $=\frac{\theta_{\star}^{(o)}\sum_{\tau=1}^{t}x_{a_{\tau}}^{2}}{\sum_{\tau=1}^{t}x_{a_{\tau}}^{2}+1}+\frac{\sum_{\tau=1}^{t}x_{a_{\tau}}g_{a_{\tau}}}{\sum_{\tau=1}^{t}x_{a_{\tau}}^{2}+1}+\frac{\sum_{\tau=1}^{t}x_{a_{\tau}}\epsilon_{\tau}}{\sum_{\tau=1}^{t}x_{a_{\tau}}^{2}+1}$
$\leq\theta_{\star}^{(o)}+\frac{g_{1}N_{1}+2g_{2}N_{2}}{N_{1}+4N_{2}+1}+\frac{\sum_{\tau=1}^{t}x_{a_{\tau}}\epsilon_{\tau}}{N_{1}+4N_{2}+1}\qquad\qquad\quad(\because\theta_{\star}^{(o)}>0)$
$\leq\theta_{\star}^{(o)}+\frac{g_{1}N_{1}+2g_{2}N_{2}}{N_{1}+4N_{2}+1}+\frac{\max_{a\in\{1,2\}}x_{a}}{N_{1}+4N_{2}+1}\sum_{\tau=1}^{t}\epsilon_{\tau}$
$=\theta_{\star}^{(o)}+\frac{g_{1}N_{1}+2g_{2}N_{2}}{N_{1}+4N_{2}+1}+\frac{2}{N_{1}+4N_{2}+1}\sum_{\tau=1}^{t}\epsilon_{\tau},$ (18)
where for the second term of Eq. 18, for $t\geq 20$, it is upper bounded by:
$\frac{g_{1}N_{1}+2g_{2}N_{2}}{N_{1}+4N_{2}+1}$ $=\frac{-3N_{1}-\frac{19}{2}N_{2}}{N_{1}+4N_{2}+1}$
$\leq\frac{-3N_{1}-\frac{19}{2}N_{2}}{N_{1}+4N_{2}+1}$
$\leq\frac{-\frac{19}{8}(N_{1}+4N_{2})}{N_{1}+4N_{2}+1}$
$=-\frac{19}{8}+\frac{19/8}{t+3N_{2}+1}$
$\leq-\frac{19}{8}+\frac{19}{8t}$
$\leq-\frac{9}{4},$
resulting in
$\begin{split}\hat{\theta}_{t}&\leq\theta_{\star}^{(o)}+\frac{g_{1}N_{1}+2g_{2}N_{2}}{N_{1}+4N_{2}+1}+\frac{2}{N_{1}+4N_{2}+1}\sum_{\tau=1}^{t}\epsilon_{\tau}\\ &\leq 2-\frac{9}{4}+\frac{2}{t+3N_{2}+1}\sum_{\tau=1}^{t}\epsilon_{\tau}\\ &=-\frac{1}{4}+\frac{2}{t+3N_{2}+1}\sum_{\tau=1}^{t}\epsilon_{\tau}\\ &\leq-\frac{1}{4}+\frac{2}{t}\sum_{\tau=1}^{t}\epsilon_{\tau}\end{split}$ (19)
Thus, we have the following:
$\mathbb{P}(\hat{\theta}_{t}>0)\leq\mathbb{P}\left(-\frac{1}{4}+\frac{2}{t}\sum_{\tau=1}^{t}\epsilon_{\tau}>0\right)=\mathbb{P}\left(\frac{2}{t}\sum_{\tau=1}^{t}\epsilon_{\tau}>\frac{1}{4}\right).$
Since $\epsilon_{\tau}$ is an IID sub-Gaussian random variable for all $\tau\in[t]$, by applying Hoeffding inequality we obtain:
$\mathbb{P}(\hat{\theta}_{t}>0)\leq\exp\left(-\frac{t}{128\sigma^{2}}\right).$
Given this, we now bound the probability of the event $E_{\widetilde{\theta}}$:
$\mathbb{P}(E_{\widetilde{\theta}})$ $=\mathbb{P}(E_{\widetilde{\theta}}\cap E_{\hat{\theta}})+\mathbb{P}(E_{\widetilde{\theta}}\cap E_{\hat{\theta}}^{c})$
$=\mathbb{P}(E_{\widetilde{\theta}}|E_{\hat{\theta}})\cdot\mathbb{P}(E_{\hat{\theta}})+\mathbb{P}(E_{\widetilde{\theta}}|E_{\hat{\theta}}^{c})\cdot\mathbb{P}(E_{\hat{\theta}}^{c})$
$=\mathbb{P}(\widetilde{\theta}_{t}\geq 0|\hat{\theta}_{t}\geq 0)\cdot\mathbb{P}(\hat{\theta}_{t}\geq 0)+\mathbb{P}(\widetilde{\theta}_{t}\geq 0|\hat{\theta}_{t}<0)\cdot\mathbb{P}(\hat{\theta}_{t}<0)$
$\leq\exp\left(-\frac{t}{128\sigma^{2}}\right)+\mathbb{P}(\widetilde{\theta}_{t}\geq 0|\hat{\theta}_{t}<0).$ (20)
Since the second term of Eq. 20 is calculated under a Gaussian distribution, which does not exceed $1/2$ for all $t\in[T]$,
$\mathbb{P}(E_{\widetilde{\theta}}^{c})\geq\frac{1}{2}-\exp\left(-\frac{t}{128\sigma^{2}}\right).$
Note that the total decision horizon $T>256\sigma^{2}\log(1/\delta^{\prime})$. For any $t>128\sigma^{2}\log(1/\delta^{\prime})$, we have $\mathbb{P}(E_{\widetilde{\theta}}^{c})\geq 1/2-\delta^{\prime}$ This implies that for the rounds more than $T/2$, the suboptimal arm is expected to be played at least $(1/2-\delta^{\prime})T/2$ times for any $\delta^{\prime}\in(0,1/2)$, thus incurring
$\mathbb{E}[\mathrm{Reg}_{\texttt{LinTS}}(T)]\geq(\gamma_{2}-\gamma_{1})\left(\frac{1}{2}-\delta^{\prime}\right)\frac{T}{2}=\frac{1}{4}\left(\frac{1}{2}-\delta^{\prime}\right)\frac{T}{2}.$
For OFUL, we also present another analysis that requires no assumption such that the suboptimal arm is played for initial $t$ rounds, which is taken in Theorem 2 of Ghosh et al. [2017]. The optimal arm, arm 2, is selected when
$x_{2}\widehat{\theta}_{t}+\frac{x_{2}}{\sqrt{1+\sum_{\tau=1}^{t-1}x_{a_{\tau}}^{2}}}>x_{1}\widehat{\theta}_{t}+\frac{x_{1}}{\sqrt{1+\sum_{\tau=1}^{t-1}x_{a_{\tau}}^{2}}},$ (21)
where $\widehat{\theta}_{t}$ is the same ridge estimator as in LinTS. The inequality Eq. 21 is equivalent to $\widehat{\theta}_{t}>(1+\sum_{\tau=1}^{t-1}x_{a_{\tau}}^{2})^{-1/2}$, which implies $\widehat{\theta}_{t}>1/\sqrt{2t}$. By Eq. 19,
$\mathbb{P}\left(\widehat{\theta}_{t}>1/\sqrt{2t}\right)$ $\leq\mathbb{P}\left(-\frac{1}{4}+\frac{2}{t}\sum_{\tau=1}^{t}\epsilon_{\tau}>\frac{1}{\sqrt{2t}}\right)$
$=\mathbb{P}\left(\frac{2}{t}\sum_{\tau=1}^{t}\epsilon_{\tau}>\frac{1}{\sqrt{2t}}+\frac{1}{4}\right)$
$\leq\mathbb{P}\left(\frac{2}{t}\sum_{\tau=1}^{t}\epsilon_{\tau}>\frac{1}{4}\right)\leq\exp\left(-\frac{t}{128\sigma^{2}}\right).$
Thus, for $t\geq 128\sigma^{2}\log(2)$, the probability of selecting arm 2 is less than $1/2$ and for $T>256\sigma^{2}\log(2)$,
$\mathbb{E}[\mathrm{Reg}_{\texttt{OFUL}}(T)]\geq(\gamma_{2}-\gamma_{1})\cdot\frac{1}{2}\cdot\frac{T}{2}=\frac{T}{16},$
and the algorithm suffers expected regret linear in $T$. ∎
C.2 Proof of Theorem 2
Let $\mathbf{V}_{t}:=\sum_{\tau=1}^{t}\sum_{a\in[K]}\mathbf{\widetilde{x}}_{a}\mathbf{\widetilde{x}}_{a}^{\top}$. Then
$\max_{a\in[K]}|{\widetilde{\mathbf{x}}_{a}}^{\top}(\widehat{\boldsymbol{\mu}}_{t}-{\boldsymbol{\mu}}_{\star})|\leq\sqrt{\sum_{a\in[K]}|{\widetilde{\mathbf{x}}_{a}}^{\top}(\widehat{\boldsymbol{\mu}}_{t}-{\boldsymbol{\mu}}_{\star})|^{2}}=t^{-1/2}\|\widehat{\boldsymbol{\mu}}_{t}-{\boldsymbol{\mu}}_{\star}\|_{\mathbf{V}_{t}}$
Recall that $\widehat{\mathbf{w}}_{t}:=\mathbf{V}_{t}^{1/2}\widehat{\boldsymbol{\mu}}_{t}$ and $\mathbf{w}_{t}:=\mathbf{V}_{t}^{1/2}\boldsymbol{\mu}_{\star}$. Then
$\max_{a\in[K]}|{\widetilde{\mathbf{x}}_{a}}^{\top}(\widehat{\boldsymbol{\mu}}_{t}-{\boldsymbol{\mu}}_{\star})|\leq t^{-1/2}\|\widehat{\mathbf{w}}_{t}-\mathbf{w}_{t}\|_{2}.$
To use Lemma 3, we prove a bound for
$\left\|\sum_{\tau=1}^{t}\sum_{a\in[K]}(\widetilde{y}_{a,\tau}-{\widetilde{\mathbf{x}}_{a}}^{\top}\mathbf{V}_{t}^{-1/2}{\mathbf{w}}_{t})\mathbf{V}_{t}^{-1/2}{\widetilde{\mathbf{x}}_{a}}\right\|_{\infty}.$
Let $\check{\mathbf{w}}_{t}^{L}:=\mathbf{V}_{t}^{1/2}\check{\boldsymbol{\mu}}_{t}^{L}$. By definition of $\widetilde{y}_{a,\tau}$,
$\left\|\sum_{\tau=1}^{t}\sum_{a\in[K]}(\widetilde{y}_{a,\tau}-\widetilde{\mathbf{x}}_{a}^{\top}\mathbf{V}_{t}^{-1/2}\mathbf{w}_{t})\mathbf{V}_{t}^{-1/2}\widetilde{\mathbf{x}}_{a}\right\|_{\infty}.$
$=\left\|\sum_{\tau=1}^{t}\sum_{a\in[K]}\left(1-\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,\tau}}\right)\mathbf{V}_{t}^{-1/2}\mathbf{\widetilde{x}}_{a}\mathbf{\widetilde{x}}_{a}^{\top}\mathbf{V}_{t}^{-1/2}\left(\check{\mathbf{w}}_{t}^{L}-\mathbf{w}_{t}\right)+\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,\tau}}\left(y_{a,\tau}-\widetilde{\mathbf{x}}_{a}^{\top}\mathbf{V}_{t}^{-1/2}\mathbf{w}_{t}\right)\mathbf{V}_{t}^{-1/2}\widetilde{\mathbf{x}}_{a}\right\|_{\infty}$
$\leq\left\|\sum_{\tau=1}^{t}\sum_{a\in[K]}\left(1-\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,\tau}}\right)\mathbf{V}_{t}^{-1/2}\mathbf{\widetilde{x}}_{a}\mathbf{\widetilde{x}}_{a}^{\top}\mathbf{V}_{t}^{-1/2}\left(\check{\mathbf{w}}_{t}^{L}-\mathbf{w}_{t}\right)\right\|_{\infty}$
$\quad+\left\|\sum_{\tau=1}^{t}\sum_{a\in[K]}\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,\tau}}\left(y_{a,\tau}-\widetilde{\mathbf{x}}_{a}^{\top}\mathbf{V}_{t}^{-1/2}\mathbf{w}_{t}\right)\mathbf{V}_{t}^{-1/2}\mathbf{\widetilde{x}}_{a}\right\|_{\infty}$
With probability at least $1-\delta$, the event $\mathcal{M}_{\tau}$ happens for all $\tau\geq 1$ and we obtain a pair of matching sample $\widetilde{a}_{\tau}$ and $a_{\tau}$. Thus, the second term is equal to,
$\left\|\sum_{\tau=1}^{t}\sum_{a\in[K]}\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,\tau}}\left(y_{a,\tau}-\widetilde{\mathbf{x}}_{a}^{\top}\mathbf{V}_{t}^{-1/2}\mathbf{w}_{t}\right)\mathbf{V}_{t}^{-1/2}\mathbf{\widetilde{x}}_{a}\right\|_{\infty}$ $=\left\|\sum_{\tau=1}^{t}\sum_{a\in[K]}\frac{\mathbb{I}(a_{\tau}=a)}{\phi_{a,\tau}}\left(y_{a,\tau}-\mathbf{\widetilde{x}}_{a}^{\top}\boldsymbol{\mu}_{\star}\right)\mathbf{V}_{t}^{-1/2}\mathbf{\widetilde{x}}_{a}\right\|_{\infty}$
$=\frac{1}{p}\left\|\sum_{\tau=1}^{t}\epsilon_{a,\tau}\mathbf{V}_{t}^{-1/2}\mathbf{\widetilde{x}}_{a_{\tau}}\right\|_{\infty}.$
Because $\|\mathbf{v}\|_{\infty}=\max_{i\in[d]}|{e_{i}^{\top}\mathbf{v}}|$ for any $\mathbf{v}\in\mathbb{R}^{d}$,
$\frac{1}{p}\left\|\sum_{\tau=1}^{t}\epsilon_{a,\tau}\mathbf{V}_{t}^{-1/2}\mathbf{\widetilde{x}}_{a_{\tau}}\right\|_{\infty}=\frac{1}{p}\max_{a\in[K]}|{\sum_{\tau=1}^{t}\epsilon_{a,\tau}\mathbf{e}_{a}^{\top}\mathbf{V}_{t}^{-1/2}\mathbf{\widetilde{x}}_{a_{\tau}}}|$
Applying Lemma 1, with probability at least $1-\delta/t^{2}$,
$\max_{a\in[K]}|{\sum_{\tau=1}^{t}\epsilon_{a,\tau}\mathbf{e}_{a}^{\top}\mathbf{V}_{t}^{-1/2}\mathbf{\widetilde{x}}_{a_{\tau}}}|$ $\leq\max_{a\in[K]}\sigma\sqrt{2\sum_{\tau=1}^{t}\left(\mathbf{e}_{a}^{\top}\mathbf{V}_{t}^{-1/2}\mathbf{\widetilde{x}}_{a_{\tau}}\right)^{2}\log\frac{2Kt^{2}}{\delta}}$
$=\max_{a\in[K]}\sigma\sqrt{2\mathbf{e}_{a}^{\top}\mathbf{V}_{t}^{-1/2}\left(\sum_{\tau=1}^{t}\mathbf{\widetilde{x}}_{a_{\tau}}\mathbf{\widetilde{x}}_{a_{\tau}}^{\top}\right)\mathbf{V}_{t}^{-1/2}\mathbf{e}_{a}\log\frac{2Kt^{2}}{\delta}}$
$\leq\max_{a\in[K]}\sigma\sqrt{2\mathbf{e}_{a}^{\top}\mathbf{V}_{t}^{-1/2}\left(\mathbf{V}_{t}\right)\mathbf{V}_{t}^{-1/2}\mathbf{e}_{a}\log\frac{2Kt^{2}}{\delta}}$
$=\sigma\sqrt{2\log\frac{2Kt^{2}}{\delta}},$
and thus,
$\frac{1}{\sqrt{p}}\left\|\sum_{\tau=1}^{t}\epsilon_{a,\tau}\mathbf{V}_{t}^{-1/2}\mathbf{\widetilde{x}}_{a_{\tau}}\right\|_{\infty}\leq\sigma\sqrt{\frac{2}{p}\log\frac{2Kt^{2}}{\delta}}$ (22)
Let $\mathbf{A}_{t}:=\sum_{\tau=1}^{t}\sum_{a\in[K]}\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,\tau}}\mathbf{\widetilde{x}}_{a}\mathbf{\widetilde{x}}_{a}^{\top}$. Then the first term,
$\begin{split}&\left\|\sum_{\tau=1}^{t}\sum_{a\in[K]}\left(1-\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,\tau}}\right)\mathbf{V}_{t}^{-1/2}\mathbf{\widetilde{x}}_{a}\mathbf{\widetilde{x}}_{a}^{\top}\mathbf{V}_{t}^{-1/2}\left(\check{\mathbf{w}}_{t}^{L}-\mathbf{w}_{t}\right)\right\|_{\infty}\\ &=\left\|\mathbf{V}_{t}^{-1/2}\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\mathbf{V}_{t}^{-1/2}\left(\check{\mathbf{w}}_{t}^{L}-\mathbf{w}_{t}\right)\right\|_{\infty}.\end{split}$ (23)
Since $\|\mathbf{v}\|_{\infty}=\max_{i\in[d]}|{\mathbf{e}_{i}^{\top}\mathbf{v}}|$ for any $\mathbf{v}\in\mathbb{R}^{d}$,
$\left\|\mathbf{V}_{t}^{-1/2}\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\mathbf{V}_{t}^{-1/2}\left(\check{\mathbf{w}}_{t}^{L}-\mathbf{w}_{t}\right)\right\|_{\infty}=$ $\max_{a\in[K]}|\mathbf{e}_{a}^{\top}\mathbf{V}_{t}^{-1/2}\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\mathbf{V}_{t}^{-1/2}\left(\check{\mathbf{w}}_{t}^{L}-\mathbf{w}_{t}\right)|$
$\leq$ $\max_{a\in[K]}\left\|\mathbf{e}_{a}^{\top}\mathbf{V}_{t}^{-1/2}\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\mathbf{V}_{t}^{-1/2}\right\|_{2}\!\!\left\|\check{\mathbf{w}}_{t}^{L}-\mathbf{w}_{t}\right\|_{2}.$
Because $\widehat{\mathbf{w}}_{t}$ is a minimizer of Eq. 8, by Lemma 3 and Eq. 22,
$\left\|\check{\mathbf{w}}_{t}^{L}-\mathbf{w}_{t}\right\|_{\mathbf{V}_{t}^{-1/2}\frac{1}{p}\sum_{\tau=1}^{t}\mathbf{\widetilde{x}}_{a_{\tau}}\mathbf{\widetilde{x}}_{a_{\tau}}^{\top}\mathbf{V}_{t}^{-1/2}}\leq 4\sigma\sqrt{\frac{2(d+d_{h})\log\frac{2Kt^{2}}{\delta}}{p\lambda_{\min}\left(\mathbf{V}_{t}^{-1/2}\frac{1}{p}\sum_{\tau=1}^{t}\mathbf{\widetilde{x}}_{a_{\tau}}\mathbf{\widetilde{x}}_{a_{\tau}}^{\top}\mathbf{V}_{t}^{-1/2}\right)}}.$
Because $\phi_{a_{\tau},\tau}=p$ and the coupling event $\cap_{\tau\geq 1}\mathcal{M}_{\tau}$ holds with probability at least $1-\delta/t^{2}$,
$\sum_{\tau=1}^{t}\frac{1}{p}\mathbf{\widetilde{x}}_{a_{\tau}}\mathbf{\widetilde{x}}_{a_{\tau}}^{\top}$ $=\sum_{\tau=1}^{t}\sum_{a\in[K]}\frac{\mathbb{I}(a_{\tau}=a)}{p}\mathbf{\widetilde{x}}_{a}\mathbf{\widetilde{x}}_{a}^{\top}$
$=\sum_{\tau=1}^{t}\sum_{a\in[K]}\frac{\mathbb{I}(a_{\tau}=a)}{\phi_{a,\tau}}\mathbf{\widetilde{x}}_{a}\mathbf{\widetilde{x}}_{a}^{\top}$
$=\sum_{\tau=1}^{t}\sum_{a\in[K]}\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,\tau}}\mathbf{\widetilde{x}}_{a}\mathbf{\widetilde{x}}_{a}^{\top}$
$:=\mathbf{A}_{t}.$
Thus, under the coupling event $\cap_{\tau=1}^{t}\mathcal{M}_{\tau}$,
$\left\|\check{\mathbf{w}}_{t}^{L}-\mathbf{w}_{t}\right\|_{\mathbf{V}_{t}^{-1/2}\mathbf{A}_{t}\mathbf{V}_{t}^{-1/2}}\leq 4\sigma\sqrt{\frac{2(d+d_{h})\log\frac{2Kt^{2}}{\delta}}{p\lambda_{\min}\left(\mathbf{V}_{t}^{-1/2}\mathbf{A}_{t}\mathbf{V}_{t}^{-1/2}\right)}}.$
By Corollary 1, with $\epsilon\in(0,1)$ to be determined later, for $t\geq 8\epsilon^{-2}(1-p)^{-2}K^{2}\log\frac{2dt^{2}}{\delta}$, with probability at least $1-\delta/t^{2}$,
$\left\|\mathbf{I}_{K}-\mathbf{V}_{t}^{-1/2}\mathbf{A}_{t}\mathbf{V}_{t}^{-1/2}\right\|_{2}\leq\epsilon,$ (24)
which implies, $(1-\epsilon)\mathbf{I}_{K}\preceq\mathbf{V}_{t}^{-1/2}\mathbf{A}_{t}\mathbf{V}_{t}^{-1/2}$ Thus,
$\left\|\check{\mathbf{w}}_{t}^{L}-\mathbf{w}_{t}\right\|_{2}\leq\frac{4\sigma}{1-\epsilon}\sqrt{\frac{2(d+d_{h})\log\frac{2Kt^{2}}{\delta}}{p}}.$
Now Eq. 23 is bounded by,
$\begin{split}&\left\|\mathbf{V}_{t}^{-1/2}\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\mathbf{V}_{t}^{-1/2}\left(\check{\mathbf{w}}_{t}^{L}-\mathbf{w}_{t}\right)\right\|_{\infty}\\ &\leq\max_{a\in[K]}\left\|\mathbf{e}_{a}^{\top}\mathbf{V}_{t}^{-1/2}\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\mathbf{V}_{t}^{-1/2}\right\|_{2}\frac{4\sigma}{1-\epsilon}\sqrt{\frac{2(d+d_{h})\log\frac{2Kt^{2}}{\delta}}{p}}.\end{split}$ (25)
With simple algebra,
$\max_{a\in[K]}\left\|\mathbf{e}_{a}^{\top}\mathbf{V}_{t}^{-1/2}\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\mathbf{V}_{t}^{-1/2}\right\|_{2}$
$=\max_{a\in[K]}\sqrt{\lambda_{\max}\left(\mathbf{V}_{t}^{-1/2}\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\mathbf{V}_{t}^{-1/2}\left(\mathbf{e}_{a}\mathbf{e}_{a}^{\top}\right)\mathbf{V}_{t}^{-1/2}\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\mathbf{V}_{t}^{-1/2}\right)}$
$\leq\max_{a\in[K]}\sqrt{\lambda_{\max}\left(\mathbf{V}_{t}^{-1/2}\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\mathbf{V}_{t}^{-1/2}\mathbf{I}_{K}\mathbf{V}_{t}^{-1/2}\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\mathbf{V}_{t}^{-1/2}\right)}$
$=\left\|\mathbf{V}_{t}^{-1/2}\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\mathbf{V}_{t}^{-1/2}\right\|_{2}$
$=\left\|\mathbf{I}_{K}-\mathbf{V}_{t}^{-1/2}\mathbf{A}_{t}\mathbf{V}_{t}^{-1/2}\right\|_{2}\leq\epsilon$
Thus,
$\left\|\mathbf{V}_{t}^{-1/2}\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\mathbf{V}_{t}^{-1/2}\left(\check{\mathbf{w}}_{t}^{L}-\mathbf{w}_{t}\right)\right\|_{\infty}\leq\frac{4\sigma\epsilon}{1-\epsilon}\sqrt{\frac{2(d+d_{h})\log\frac{2Kt^{2}}{\delta}}{p}}.$
Now we obtain,
$\left\|\sum_{\tau=1}^{t}\sum_{a\in[K]}(\widetilde{y}_{a,\tau}-\widetilde{\mathbf{x}}_{a}^{\top}\mathbf{V}_{t}^{-1/2}\mathbf{w}_{t})\mathbf{V}_{t}^{-1/2}\widetilde{\mathbf{x}}_{a}\right\|_{\infty}$ $\leq\frac{4\sigma\epsilon}{1-\epsilon}\sqrt{\frac{2(d+d_{h})\log\frac{2Kt^{2}}{\delta}}{p}}+\frac{\sigma}{p}\sqrt{2\log\frac{2Kt^{2}}{\delta}}$
$\leq\frac{4\sigma\epsilon}{1-\epsilon}\sqrt{\frac{2K\log\frac{2Kt^{2}}{\delta}}{p}}+\frac{\sigma}{p}\sqrt{2\log\frac{2Kt^{2}}{\delta}}$
$=\left(\frac{4\epsilon\sqrt{K}}{1-\epsilon}+\frac{1}{p}\right)\sigma\sqrt{2\log\frac{2Kt^{2}}{\delta}}$
Setting $\epsilon=p^{-1}/(\sqrt{K}+p^{-1})$ gives $\frac{\epsilon\sqrt{K}}{{1-\epsilon}}=p^{-1}$ and for $t\geq 8(\sqrt{K}+p^{-1})^{2}p^{2}(1-p)^{-2}K^{2}\log\frac{2Kt^{2}}{\delta}$,
$\left\|\sum_{\tau=1}^{t}\sum_{a\in[K]}(\widetilde{y}_{a,\tau}-\widetilde{\mathbf{x}}_{a}^{\top}\mathbf{V}_{t}^{-1/2}\mathbf{w}_{t})\mathbf{V}_{t}^{-1/2}\widetilde{\mathbf{x}}_{a}\right\|_{\infty}\leq\frac{5\sigma}{p}\sqrt{2\log\frac{2Kt^{2}}{\delta}}$
Because $\widehat{\mathbf{w}}_{t}$ is a minimizer of (11), by Lemma 3,
$\left\|\widehat{\mathbf{w}}_{t}-\mathbf{w}_{t}\right\|_{\mathbf{V}_{t}^{-1/2}(\sum_{\tau=1}^{t}\widetilde{\mathbf{x}}_{a}\widetilde{\mathbf{x}}_{a})\mathbf{V}_{t}^{-1/2}}\leq\frac{20\sigma}{p}\sqrt{\frac{2(d+d_{h})\log\frac{2Kt^{2}}{\delta}}{\lambda_{\min}\left(\mathbf{V}_{t}^{-1/2}(\sum_{\tau=1}^{t}\widetilde{\mathbf{x}}_{a}\widetilde{\mathbf{x}}_{a})\mathbf{V}_{t}^{-1/2}\right)}},$
which is equivalent to,
$\left\|\widehat{\mathbf{w}}_{t}-\mathbf{w}_{t}\right\|_{2}\leq\frac{20\sigma}{p}\sqrt{2(d+d_{h})\log\frac{2Kt^{2}}{\delta}}.$
This concludes,
$\max_{a\in[K]}|\widetilde{\mathbf{x}}_{a}(\widehat{\boldsymbol{\mu}}_{t}-\boldsymbol{\mu}_{\star})|\leq\frac{20\sigma}{p}\sqrt{\frac{2(d+d_{h})\log\frac{2Kt^{2}}{\delta}}{t}},$
which conmpletes the proof. ∎
C.3 Proof of Theorem 3
Because the regret is bounded by $2$ and the number of rounds for the exploration phase is at most $|\mathcal{E}_{T}|\leq 8(\sqrt{K}+p^{-1})^{2}p^{2}(1-p)^{-2}K^{2}\log\frac{2KT^{2}}{\delta}$.
$\mathrm{Reg}(T)$ $\leq\frac{16K^{2}(\sqrt{K}+p^{-1})^{2}}{(1-p)^{2}}\log\frac{2Kt^{2}}{\delta}+\sum_{t\in[T]\setminus\mathcal{E}_{T}}\mathbb{E}_{t-1}[y_{\star,t}]-\mathbb{E}_{t-1}[y_{a_{t},t}]$
$=$ $\frac{16K^{2}(\sqrt{K}+p^{-1})^{2}}{(1-p)^{2}}+\sum_{t\in[T]\setminus\mathcal{E}_{T}}\left\{\mathbb{I}\left(a_{t}=\widehat{a}_{t}\right)\left(\mathbb{E}_{t-1}[y_{\star,t}]-\mathbb{E}_{t-1}[y_{a_{t},t}]\right)\right\}$
$+\sum_{t\in[T]\setminus\mathcal{E}_{T}}\left\{\mathbb{I}\left(a_{t}\neq\widehat{a}_{t}\right)\left(\mathbb{E}_{t-1}[y_{\star,t}]-\mathbb{E}_{t-1}[y_{a_{t},t}]\right)\right\}.$
By Theorem 2, on the event $\{a_{t}=\widehat{a}_{t}\}$,
$\mathbb{E}_{t-1}[y_{\star,t}]-\mathbb{E}_{t-1}[y_{a_{t},t}]$ $=\widetilde{\boldsymbol{x}}_{a_{\star}}^{\top}{\boldsymbol{\mu}}_{\star}-\widetilde{\boldsymbol{x}}_{\widehat{a}_{t}}^{\top}{\boldsymbol{\mu}}_{\star}$
$\leq 2\max_{a\in[K]}\left|\widetilde{\mathbf{x}}_{a}^{\top}\left({\boldsymbol{\mu}}_{\star}-\widehat{\boldsymbol{\mu}}_{t-1}^{L}\right)\right|+\widetilde{\boldsymbol{x}}_{a_{\star}}^{\top}\widehat{\boldsymbol{\mu}}_{t-1}^{L}-\widetilde{\boldsymbol{x}}_{\widehat{a}_{t}}^{\top}\widehat{\boldsymbol{\mu}}_{t-1}^{L}$
$\leq 2\max_{a\in[K]}\left|\widetilde{\mathbf{x}}_{a}^{\top}\left({\boldsymbol{\mu}}_{\star}-\widehat{\boldsymbol{\mu}}_{t-1}^{L}\right)\right|$
$\leq\frac{40\sigma}{p}\sqrt{\frac{2(d+d_{h})}{t}\log\frac{2Kt^{2}}{\delta}},$
with probability at least $1-5\delta/t^{2}$ for each $t\in[T]\setminus\mathcal{E}_{T}$. Summing over $t$ gives,
$\sum_{t\in[T]\setminus\mathcal{E}_{T}}\left\{\mathbb{I}\left(a_{t}=\widehat{a}_{t}\right)\left(\mathbb{E}_{t-1}[y_{\star,t}]-\mathbb{E}_{t-1}[y_{a_{t},t}]\right)\right\}\leq\frac{80\sigma}{p}\sqrt{\frac{2(d+d_{h})}{t}\log\frac{2Kt^{2}}{\delta}}.$
By resampling at most $\rho_{t}$ times, the probability of the event $\{a_{t}\neq\widehat{a}_{t}\}$ is
$\mathbb{P}\left(a_{t}\neq\widehat{a}_{t}\right)=$ $\sum_{m=1}^{\rho_{t}}\frac{p}{(K-1)\sqrt{t}}\left(1-\frac{p}{(K-1)\sqrt{t}}\right)^{m-1}\cdot\mathbb{P}(\text{Resample succeed at trial }m)$
$\leq$ $\sum_{m=1}^{\rho_{t}}\frac{p}{(K-1)\sqrt{t}}\left(1-\frac{p}{(K-1)\sqrt{t}}\right)^{m-1}$
$=$ $\frac{p}{(K-1)\sqrt{t}}\left(\frac{p}{(K-1)\sqrt{t}}\right)^{-1}\left\{1-\left(1-\frac{p}{(K-1)\sqrt{t}}\right)^{\rho_{t}}\right\}$
$=$ $1-\left(1-\frac{p}{(K-1)\sqrt{t}}\right)^{\rho_{t}}$
$\leq$ $\frac{p\rho_{t}}{(K-1)\sqrt{t}},$
where the last inequality uses $(1+x)^{n}\geq 1+nx$ for $x\geq-1$ and $n\in\mathbb{N}$. Then the expected sum of regret,
$\mathbb{E}\left[\sum_{t\in[T]\setminus\mathcal{E}_{T}}\left\{\mathbb{I}\left(a_{t}=\widehat{a}_{t}\right)\left(\mathbb{E}_{t-1}[y_{\star,t}]-\mathbb{E}_{t-1}[y_{a_{t},t}]\right)\right\}\right]$
$\leq 2\sum_{t\in[T]\setminus\mathcal{E}_{T}}\mathbb{P}\left(a_{t}\neq\widehat{a}_{t}\right)$
$\leq\frac{4p\sqrt{T}}{K-1}\rho_{T}$
$=\frac{4p\sqrt{T}}{K-1}\frac{\log\frac{(T+1)^{2}}{\delta}}{\log(1/p)}.$
Thus,
$\mathbb{E}[\mathrm{Reg}(T)]$ $\leq 10\delta\log T+\frac{16K^{2}(\sqrt{K}+p^{-1})^{2}}{(1-p)^{2}}+\frac{4p\sqrt{T}}{K-1}\frac{\log\frac{(T+1)^{2}}{\delta}}{\log(1/p)}$
$\quad+\frac{80\sigma}{p}\sqrt{2(d+d_{h})T\log\frac{2KT^{2}}{\delta}},$
which concludes the proof. ∎
C.4 Proof of Theorem 4
Let $\widetilde{\mathbf{V}}_{t}:=\sum_{\tau=1}^{t}\mathbb{I}(\mathcal{M}_{\tau})\sum_{a\in[K]}\mathbf{\widetilde{x}}_{a}\mathbf{\widetilde{x}}_{a}^{\top}+\mathbf{I}_{K}$ and $\mathbf{V}_{t}:=\sum_{\tau=1}^{t}\sum_{a\in[K]}\mathbf{\widetilde{x}}_{a}\mathbf{\widetilde{x}}_{a}^{\top}+\mathbf{I}_{K}$. By definition of $\widehat{\boldsymbol{\mu}}_{t}^{R}$,
$\widetilde{\mathbf{x}}_{a}^{\top}(\widehat{\boldsymbol{\mu}}_{t}-{\boldsymbol{\mu}}_{\star})=\widetilde{\mathbf{x}}_{a}^{\top}\widetilde{\mathbf{V}}_{t}^{-1}\left\{\sum_{\tau=1}^{t}\mathbb{I}(\mathcal{M}_{\tau})\sum_{a\in[K]}\widetilde{\mathbf{x}}_{a}\left(\widetilde{y}_{a,\tau}-\widetilde{\mathbf{x}}_{a}^{\top}{\boldsymbol{\mu}}_{\star}\right)-{\boldsymbol{\mu}}_{\star}\right\}.$
By definition of the pseudo-rewards,
$\widetilde{y}_{a,\tau}-\widetilde{\mathbf{x}}_{a}^{\top}{\boldsymbol{\mu}}_{\star}=\left(1-\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,t}}\right)\widetilde{\mathbf{x}}_{a}^{\top}\left(\check{\boldsymbol{\mu}}_{t}^{R}-{\boldsymbol{\mu}}_{\star}\right)+\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,\tau}}\epsilon_{a,\tau}.$
Let $\widetilde{\mathbf{A}}_{t}:=\sum_{\tau=1}^{t}\mathbb{I}(\mathcal{M}_{\tau})\sum_{a\in[K]}\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,t}}\mathbf{\widetilde{x}}_{a}\mathbf{\widetilde{x}}_{a}^{\top}+\mathbf{I}_{K}$ and $\mathbf{A}_{t}:=\sum_{\tau=1}^{t}\sum_{a\in[K]}\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,t}}\mathbf{\widetilde{x}}_{a}\mathbf{\widetilde{x}}_{a}^{\top}+\mathbf{I}_{K}$ Then,
$\widetilde{\mathbf{x}}_{a}^{\top}(\widehat{\boldsymbol{\mu}}_{t}-{\boldsymbol{\mu}}_{\star})=\widetilde{\mathbf{x}}_{a}^{\top}\widetilde{\mathbf{V}}_{t}^{-1}\left\{\left(\widetilde{\mathbf{V}}_{t}-\widetilde{\mathbf{A}}_{t}\right)\left(\check{\boldsymbol{\mu}}_{t}^{R}-{\boldsymbol{\mu}}_{\star}\right)+\sum_{\tau=1}^{t}\mathbb{I}(\mathcal{M}_{\tau})\sum_{a\in[K]}\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,\tau}}\widetilde{\mathbf{x}}_{a}\epsilon_{a,\tau}-{\boldsymbol{\mu}}_{\star}\right\}.$
By definition of the imputation estimator $\check{\boldsymbol{\mu}}_{t}$,
$\check{\boldsymbol{\mu}}_{t}^{R}-{\boldsymbol{\mu}}_{\star}=$ $\left(\sum_{\tau=1}^{t}\widetilde{\boldsymbol{x}}_{a_{\tau}}\widetilde{\boldsymbol{x}}_{a_{\tau}}^{\top}+p\mathbf{I}_{K}\right)^{-1}\left(\sum_{\tau=1}^{t}\widetilde{\boldsymbol{x}}_{a_{\tau}}\epsilon_{a_{\tau},\tau}-p{\boldsymbol{\mu}}_{\star}\right)$
$=$ $\left(\sum_{\tau=1}^{t}\frac{1}{\phi_{a_{\tau},\tau}}\widetilde{\boldsymbol{x}}_{a_{\tau}}\widetilde{\boldsymbol{x}}_{a_{\tau}}^{\top}+\mathbf{I}_{K}\right)^{-1}\left(\sum_{\tau=1}^{t}\frac{1}{\phi_{a_{\tau},\tau}}\widetilde{\boldsymbol{x}}_{a_{\tau}}\epsilon_{a_{\tau},\tau}-{\boldsymbol{\mu}}_{\star}\right)$
$=\left(\sum_{\tau=1}^{t}\sum_{a\in[K]}\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,\tau}}\widetilde{\boldsymbol{x}}_{a_{\tau}}\widetilde{\boldsymbol{x}}_{a_{\tau}}^{\top}+\mathbf{I}_{K}\right)^{-1}\left(\sum_{\tau=1}^{t}\frac{1}{p}\widetilde{\boldsymbol{x}}_{a_{\tau}}\epsilon_{a_{\tau},\tau}-{\boldsymbol{\mu}}_{\star}\right),$
where the second equality holds because $\phi_{a_{\tau},\tau}=p$. Under the coupling event $\cap_{\tau=1}^{t}\mathcal{M}_{\tau}$,
$\sum_{\tau=1}^{t}\mathbb{I}(\mathcal{M}_{\tau})\sum_{a\in[K]}\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,t}}\widetilde{\mathbf{x}}_{a}\widetilde{\mathbf{x}}_{a}^{\top}+\mathbf{I}_{K}=$ $\sum_{\tau=1}^{t}\sum_{a\in[K]}\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,t}}\widetilde{\mathbf{x}}_{a}\widetilde{\mathbf{x}}_{a}^{\top}+\mathbf{I}_{K}$
::=:= $\mathbf{A}_{t},$
and
$\widetilde{\mathbf{x}}_{a}^{\top}(\widehat{\boldsymbol{\mu}}_{t}-{\boldsymbol{\mu}}_{\star})=$ $\widetilde{\mathbf{x}}_{a}^{\top}\mathbf{V}_{t}^{-1}\left\{\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\mathbf{A}_{t}^{-1}\left(\sum_{\tau=1}^{t}\frac{1}{p}\widetilde{\boldsymbol{x}}_{a_{\tau}}\epsilon_{a_{\tau},\tau}-{\boldsymbol{\mu}}_{\star}\right)+\sum_{\tau=1}^{t}\frac{\epsilon_{a_{\tau}}}{\phi_{a_{\tau},\tau}}\widetilde{\boldsymbol{x}}_{a_{\tau}}-{\boldsymbol{\mu}}_{\star}\right\}$
$=$ $\widetilde{\mathbf{x}}_{a}^{\top}\mathbf{V}_{t}^{-1}\left\{\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\mathbf{A}_{t}^{-1}+\mathbf{I}_{K}\right\}\left(\sum_{\tau=1}^{t}\frac{1}{p}\widetilde{\boldsymbol{x}}_{a_{\tau}}\epsilon_{a_{\tau},\tau}-{\boldsymbol{\mu}}_{\star}\right)$
$=$ $\widetilde{\mathbf{x}}_{a}^{\top}\mathbf{V}_{t}^{-1/2}\left(\mathbf{V}_{t}^{1/2}\mathbf{A}_{t}^{-1}\mathbf{V}_{t}^{1/2}\right)\mathbf{V}_{t}^{-1/2}\left(\sum_{\tau=1}^{t}\frac{1}{p}\widetilde{\boldsymbol{x}}_{a_{\tau}}\epsilon_{a_{\tau},\tau}-{\boldsymbol{\mu}}_{\star}\right).$
Taking absolute value on both sides, by Cauchy-Schwarz inequality,
$\max_{a\in[K]}|\widetilde{\mathbf{x}}_{a}^{\top}(\widehat{\boldsymbol{\mu}}_{t}-{\boldsymbol{\mu}}_{\star})|\leq\max_{a\in[K]}\|\widetilde{\mathbf{x}}_{a}\|_{\mathbf{V}_{t}^{-1}}\|\mathbf{V}_{t}^{1/2}\mathbf{A}_{t}^{-1}\mathbf{V}_{t}^{1/2}\|_{2}\left\|\sum_{\tau=1}^{t}\frac{1}{p}\widetilde{\boldsymbol{x}}_{a_{\tau}}\epsilon_{a_{\tau},\tau}-{\boldsymbol{\mu}}_{\star}\right\|_{\mathbf{V}_{t}^{-1}}.$
By Corollary which implies $\mathbf{I}_{K}-\mathbf{V}_{t}^{-1/2}\mathbf{A}_{t}\mathbf{V}_{t}^{-1/2}\preceq\epsilon\mathbf{I}_{K}.$ Rearraging the terms,
$\mathbf{V}_{t}^{1/2}\mathbf{A}_{t}^{-1}\mathbf{V}_{t}^{1/2}\preceq(1-\epsilon)^{-1}\mathbf{I}_{K}.$
Thus,
$\max_{a\in[K]}|\widetilde{\mathbf{x}}_{a}^{\top}(\widehat{\boldsymbol{\mu}}_{t}-{\boldsymbol{\mu}}_{\star})|$ $\leq\frac{\max_{a\in[K]}\|\widetilde{\mathbf{x}}_{a}\|_{\mathbf{V}_{t}^{-1}}}{1-\epsilon}\left\|\sum_{\tau=1}^{t}\frac{1}{p}\widetilde{\boldsymbol{x}}_{a_{\tau}}\epsilon_{a_{\tau},\tau}-{\boldsymbol{\mu}}_{\star}\right\|_{\mathbf{V}_{t}^{-1}}$
$\leq\frac{\max_{a\in[K]}\|\widetilde{\mathbf{x}}_{a}\|_{\mathbf{V}_{t}^{-1}}}{1-\epsilon}\left(\frac{1}{p}\left\|\sum_{\tau=1}^{t}\widetilde{\boldsymbol{x}}_{a_{\tau}}\epsilon_{a_{\tau},\tau}\right\|_{\mathbf{V}_{t}^{-1}}+\left\|{\boldsymbol{\mu}}_{\star}\right\|_{\mathbf{V}_{t}^{-1}}\right).$
Note that the matrix $\mathbf{V}_{t}$ is deterministic. By Lemma 9 in [Abbasi-yadkori et al., 2011], with probability at least $1-\delta,$
$\left\|\sum_{\tau=1}^{t}\widetilde{\boldsymbol{x}}_{a_{\tau}}\epsilon_{a_{\tau},\tau}\right\|_{\mathbf{V}_{t}^{-1}}$ $\leq\left\|\sum_{\tau=1}^{t}\widetilde{\boldsymbol{x}}_{a_{\tau}}\epsilon_{a_{\tau},\tau}\right\|_{\left(\sum_{\tau=1}^{t}\widetilde{\boldsymbol{x}}_{a_{\tau}}\widetilde{\boldsymbol{x}}_{a_{\tau}}^{\top}+\mathbf{I}_{K}\right)^{-1}}$
$\leq\sigma\sqrt{2\log\frac{\det(\sum_{\tau=1}^{t}\widetilde{\boldsymbol{x}}_{a_{\tau}}\widetilde{\boldsymbol{x}}_{a_{\tau}}^{\top}+\mathbf{I}_{K}){}^{1/2}}{\delta}}$
$\leq\sigma\sqrt{\log\frac{\det(\sum_{\tau=1}^{t}\widetilde{\boldsymbol{x}}_{a_{\tau}}\widetilde{\boldsymbol{x}}_{a_{\tau}}^{\top}+\mathbf{I}_{K})}{\delta}},$
for all $t\geq 1$. Because
$\det\left(\sum_{\tau=1}^{t}\widetilde{\boldsymbol{x}}_{a_{\tau}}\widetilde{\boldsymbol{x}}_{a_{\tau}}^{\top}+\mathbf{I}_{K}\right)$$\leq\left\{\frac{\text{Tr}\left(\sum_{\tau=1}^{t}\widetilde{\boldsymbol{x}}_{a_{\tau}}\widetilde{\boldsymbol{x}}_{a_{\tau}}^{\top}\right)+K}{K}\right\}^{K}$
$\leq\left\{\frac{t\max_{a\in[K]}\|\widetilde{\boldsymbol{x}}_{a_{\tau}}\|_{2}+K}{K}\right\}^{K}$
$\leq\left\{t+1\right\}^{K},$
where the last inequality holds by $\|\widetilde{\boldsymbol{x}}_{a_{\tau}}\|_{2}\leq\sqrt{K}\|\widetilde{\boldsymbol{x}}_{a_{\tau}}\|_{\infty}\leq K$. Thus,
$\left\|\sum_{\tau=1}^{t}\widetilde{\boldsymbol{x}}_{a_{\tau}}\epsilon_{a_{\tau},\tau}\right\|_{\mathbf{V}_{t}^{-1}}\leq\sigma\sqrt{K\log\frac{t+1}{\delta}},$
which proves,
$\max_{a\in[K]}|\widetilde{\mathbf{x}}_{a}^{\top}(\widehat{\boldsymbol{\mu}}_{t}-{\boldsymbol{\mu}}_{\star})|$ $\leq\frac{\max_{a\in[K]}\|\widetilde{\mathbf{x}}_{a}\|_{\mathbf{V}_{t}^{-1}}}{1-\epsilon}\left(\frac{\sigma}{p}\sqrt{K\log\frac{t+1}{\delta}}+\left\|{\boldsymbol{\mu}}_{\star}\right\|_{\mathbf{V}_{t}^{-1}}\right)$
$\leq\frac{1}{\sqrt{t}}\cdot\frac{1}{1-\epsilon}\left(\frac{\sigma}{p}\sqrt{K\log\frac{t+1}{\delta}}+\left\|{\boldsymbol{\mu}}_{\star}\right\|_{\mathbf{V}_{t}^{-1}}\right).$
Because $\|{\boldsymbol{\mu}}_{\star}\|_{\mathbf{V}_{t}^{-1}}\leq\|{\boldsymbol{\mu}}_{\star}\|_{2}\leq\sqrt{K}$, setting $\epsilon=1/2$ completes the proof. ∎
C.5 Proof of Theorem 5
Because the regret is bounded by $1$ and the number of rounds for the exploration phase is at most $|\mathcal{E}_{T}|\leq 32(1-p)^{-2}K^{2}\log\frac{2dT^{2}}{\delta}.$
$\mathrm{Reg}(T)$ $\leq 32(1-p)^{-2}K^{2}\log\frac{2dT^{2}}{\delta}+\sum_{t\in[T]\setminus\mathcal{E}_{T}}\mathbb{E}_{t-1}[y_{\star,t}]-\mathbb{E}_{t-1}[y_{a_{t},t}]$
$=$ $32(1-p)^{-2}K^{2}\log\frac{2dT^{2}}{\delta}+\sum_{t\in[T]\setminus\mathcal{E}_{T}}\left\{\mathbb{I}\left(a_{t}=\widehat{a}_{t}\right)\left(\mathbb{E}_{t-1}[y_{\star,t}]-\mathbb{E}_{t-1}[y_{a_{t},t}]\right)\right\}$
$+\sum_{t\in[T]\setminus\mathcal{E}_{T}}\left\{\mathbb{I}\left(a_{t}\neq\widehat{a}_{t}\right)\left(\mathbb{E}_{t-1}[y_{\star,t}]-\mathbb{E}_{t-1}[y_{a_{t},t}]\right)\right\}.$
On the event $\{a_{t}=\widehat{a}_{t}\}$,
$\mathbb{E}_{t-1}[y_{\star,t}]-\mathbb{E}_{t-1}[y_{a_{t},t}]=$ $\widetilde{\boldsymbol{x}}_{a_{\star}}^{\top}{\boldsymbol{\mu}}_{\star}-\widetilde{\boldsymbol{x}}_{\widehat{a}_{t}}^{\top}{\boldsymbol{\mu}}_{\star}$
$\leq 2\max_{a\in[K]}\left|\widetilde{\mathbf{x}}_{a}^{\top}\left({\boldsymbol{\mu}}_{\star}-\widehat{\boldsymbol{\mu}}_{t-1}^{R}\right)\right|+\widetilde{\boldsymbol{x}}_{a_{\star}}^{\top}\widehat{\boldsymbol{\mu}}_{t-1}^{R}-\widetilde{\boldsymbol{x}}_{\widehat{a}_{t}}^{\top}\widehat{\boldsymbol{\mu}}_{t-1}^{R}$
$\leq 2\max_{a\in[K]}\left|\widetilde{\mathbf{x}}_{a}^{\top}\left({\boldsymbol{\mu}}_{\star}-\widehat{\boldsymbol{\mu}}_{t-1}^{R}\right)\right|$
$\leq\frac{4}{\sqrt{t}}\left(\frac{\sigma}{p}\sqrt{K\log\frac{2t^{2}}{\delta}}+\sqrt{K}\right),$
with probability at least $1-3\delta/t$, by Theorem 4. Summing over $t$ gives,
$\sum_{t\in[T]\setminus\mathcal{E}_{T}}\left\{\mathbb{I}\left(a_{t}=\widehat{a}_{t}\right)\left(\mathbb{E}_{t-1}[y_{\star,t}]-\mathbb{E}_{t-1}[y_{a_{t},t}]\right)\right\}\leq 8\sqrt{KT}\left(\frac{\sigma}{p}\sqrt{\log\frac{2T^{2}}{\delta}}+1\right).$
By resampling at most $\rho_{t}$ times, the probability of the event $\{a_{t}\neq\widehat{a}_{t}\}$ is
$\mathbb{P}\left(a_{t}\neq\widehat{a}_{t}\right)=$ $\sum_{m=1}^{\rho_{t}}\frac{p}{(K-1)\sqrt{t}}\left(1-\frac{p}{(K-1)\sqrt{t}}\right)^{m-1}$
$=$ $\frac{p}{(K-1)\sqrt{t}}\left(\frac{p}{(K-1)\sqrt{t}}\right)^{-1}\left\{1-\left(1-\frac{p}{(K-1)\sqrt{t}}\right)^{\rho_{t}}\right\}$
$=$ $1-\left(1-\frac{p}{(K-1)\sqrt{t}}\right)^{\rho_{t}}$
$\geq$ $\frac{p\rho_{t}}{(K-1)\sqrt{t}},$
where the last inequality uses $(1+x)^{n}\geq 1+nx$ for $x\geq-1$ and $n\in\mathbb{N}$. Then the expected sum of regret,
$\mathbb{E}\left[\sum_{t\in[T]\setminus\mathcal{E}_{T}}\left\{\mathbb{I}\left(a_{t}=\widehat{a}_{t}\right)\left(\mathbb{E}_{t-1}[y_{\star,t}]-\mathbb{E}_{t-1}[y_{a_{t},t}]\right)\right\}\right]$ $\leq\sum_{t\in[T]\setminus\mathcal{E}_{T}}\mathbb{P}\left(a_{t}\neq\widehat{a}_{t}\right)$
$\leq\frac{2p\sqrt{T}}{K-1}\rho_{T}$
$=\frac{2p\sqrt{T}}{K-1}\frac{\log\frac{(T+1)^{2}}{\delta}}{\log(1/p)}.$
Thus,
$\mathbb{E}[\mathrm{Reg}(T)]\leq 6\delta\log T+\frac{32K^{2}}{(1-p)^{2}}\log\frac{2dT^{2}}{\delta}+\frac{2p\sqrt{T}}{K-1}\frac{\log\frac{(T+1)^{2}}{\delta}}{\log(1/p)}+8\sqrt{KT}\left(\frac{\sigma}{p}\sqrt{\log\frac{T+1}{\delta}}+1\right).\hfill\qed$
Appendix D Algorithm-agnostic Lower Bound of Regret Ignoring Unobserved Features
In this section, extending our argument in Theorem 1, we show that there exists a problem instance where linear bandit algorithms relying solely on observed features can incur regret that grows linearly in $T$. We begin by formally defining such algorithms.
Definition 1 (Policy dependent on observed features).
For each $t\in[T]$, let $\pi_{t}:\mathbb{R}^{d}\times\mathbb{R}^{t-1}\to[0,1]$, be a policy that maps an observed feature vector $\mathbf{x}\in\{\mathbf{x}_{a}:a\in[K]\}$, given past reward observation $\{y_{a_{s},s}:s\in[t-1]\}$, to a probability of selection. Then the policy $\pi_{t}$ is dependent only on observed features if, for any $y_{a_{1},1},\ldots,y_{a_{t-1},t-1}$, it holds that $\mathbf{x}_{1}=\mathbf{x}_{2}$ implies $\pi_{t}(\mathbf{x}_{1}|y_{a_{1},1},\ldots,y_{a_{t-1},t-1})=\pi_{t}(\mathbf{x}_{2}|y_{a_{1},1},\ldots,y_{a_{t-1},t-1})$.
For instance, the UCB and Thompson sampling-based policies for linear bandits (with observed features), considered in Theorem 1, satisfy Definition 1, as they assign the same selection probability as long as the observed features are the same. In contrast, the policy in the MAB algorithms (that disregard observed features) may assign different selection probability although the observed features are equal and is not dependent on the observed features. In the theorem below, we particularly provide a lower bound for algorithms that employ policies that are dependent on the observed features.
Theorem 7 (Regret Lower Bound under Policies Dependent on Observed Features).
For any algorithm $\Pi:=(\pi_{1},\ldots,\pi_{T})$ that consists of policies $\{\pi_{t}:t\in[T]\}$ that are dependent on observed features, there exists a set of features $\{\mathbf{z}_{1},\ldots,\mathbf{z}_{K}\}$ and a parameter $\boldsymbol{\theta}_{\star}\in\mathbb{R}^{d_{z}}$ such that the cumulative regret
$\mathrm{Reg}_{\Pi}(T,\theta_{\star},\mathbf{z}_{1},\ldots,\mathbf{z}_{K})\geq\frac{T}{6}.$
Proof.
We start the proof by providing a detailed account of the scenario described in the theorem. Without loss of generality, we consider the case where $K=3$. As stated in the theorem, $a_{\star}$ represents the index of the optimal action when considering the entire reward, including both observed and latent components. In contrast, $a_{o}$ denotes the index of the optimal action when considering only the observed components. We introduce an additional notation, $a^{\prime}$, which refers to an action whose observed features are identical to those of $a_{\star}$, but with a distinct latent component. Specifically, this implies that $a^{\prime}\neq a_{\star}$ and $\mathbf{z}_{a^{\prime}}\neq\mathbf{z}_{a_{\star}}$, but $\mathbf{x}_{a^{\prime}}=\mathbf{x}_{a_{\star}}$. By definition of the policy $\pi_{t}$ that depends on the observed features, $\pi_{t}(\mathbf{x}_{a_{\star}})=\pi_{t}(\mathbf{x}_{a^{\prime}})$ and the probability of selecting an optimal arm is $\pi_{t}(\mathbf{x}_{a_{\star}})\leq 1/2$.
Taking this scenario into account, the observed part of the features associated with $a_{\star}$, $a^{\prime}$, and $a_{o}$ are defined as follows:
$\mathbf{x}_{a_{\star}}:=\left[-\frac{1}{2},\dots,-\frac{1}{2}\right]^{\top},\mathbf{x}_{a^{\prime}}:=\left[-\frac{1}{2},\dots,-\frac{1}{2}\right]^{\top},\mathbf{x}_{a_{o}}:=\left[\frac{1}{2},\dots,\frac{1}{2}\right]^{\top}.$
Additionally, we define the unobserved feature vectors for actions $a_{\star}$, $a^{\prime}$, and $a_{o}$ as follows:
$\mathbf{u}_{a_{\star}}:=\left[1,\dots,1\right]^{\top},\mathbf{u}_{a^{\prime}}:=\left[-1,\dots,-1\right]^{\top},\mathbf{u}_{a_{o}}:=\left[-1,\dots,-1,1,\dots,1\right]^{\top},$
where in $\mathbf{u}_{a_{o}}$, the number of 1’s and -1’s are equal. This ensures that the scenario aligns with the assumption imposed on the feature vectors throughout this paper. We further define the true parameter as follows:
$\boldsymbol{\theta}_{\star}:=\left[{\frac{1}{3d}},\dots,{\frac{1}{3d}},{\frac{2}{3d_{u}}},\dots,{\frac{2}{3d_{u}}}\right]^{\top}\in\mathbb{R}^{d_{z}},$
thus it follows that $\boldsymbol{\theta}_{\star}^{(o)}=[1/3d,\dots,1/3d]^{\top}\in\mathbb{R}^{d}$ and $\boldsymbol{\theta}_{\star}^{(u)}=[2/3d_{u},\dots,2/3d_{u}]^{\top}\in\mathbb{R}^{d_{u}}$. Note that it is straightforward to verify that $|\langle\mathbf{z}_{a},\boldsymbol{\theta}_{\star}\rangle|\leq 1$, thereby satisfying the assumption on the mean reward (Section 3.2). With this established, we can also observe that the expected reward for the three actions are defined as:
$\langle{\mathbf{z}_{a_{\star}}},{\boldsymbol{\theta}_{\star}}\rangle=\langle{\mathbf{x}_{a_{\star}}},{\boldsymbol{\theta}_{\star}^{(o)}}\rangle+\langle{\mathbf{u}_{a_{\star}}},{\boldsymbol{\theta}_{\star}^{(u)}}\rangle=-\frac{1}{6}+\frac{2}{3}=\frac{1}{2},$
$\langle{{\mathbf{z}}_{a^{\prime}}},{\boldsymbol{\theta}_{\star}}\rangle=\langle{\mathbf{x}_{a^{\prime}}},{\boldsymbol{\theta}_{\star}^{(o)}}\rangle+\langle{\mathbf{u}_{a^{\prime}}},{\boldsymbol{\theta}_{\star}^{(u)}}\rangle=-\frac{1}{6}-\frac{2}{3}=-\frac{5}{6},$
$\langle{{\mathbf{z}}_{a_{o}}},{\boldsymbol{\theta}_{\star}}\rangle=\langle{\mathbf{x}_{a_{o}}},{\boldsymbol{\theta}_{\star}^{(o)}}\rangle+\langle{\mathbf{u}_{a_{o}}},{\boldsymbol{\theta}_{\star}^{(u)}}\rangle=\frac{1}{6}+0=\frac{1}{6},$
respectively, and it is straightforward to verify that $\langle{\mathbf{z}_{a_{\star}}},{\boldsymbol{\theta}_{\star}}\rangle-\langle{{\mathbf{z}}_{a_{o}}},{\boldsymbol{\theta}_{\star}}\rangle=2/3>0$ and that $\langle{\mathbf{z}_{a_{\star}}},{\boldsymbol{\theta}_{\star}}\rangle-\langle{{\mathbf{z}}_{a^{\prime}}},{\boldsymbol{\theta}_{\star}}\rangle=4/3>0$, which confirms that $a_{\star}$ is optimal when considering the full feature set.
At each round $t\in[T]$, for any policy $\pi_{t}$ satisfying Definition 1, we have $\pi_{t}(\mathbf{x}_{a_{\star}}|y_{a_{1},1},\ldots,y_{a_{t-1},t-1})=\pi_{t}(\mathbf{x}_{a^{\prime}}|y_{a_{1},1},\ldots,y_{a_{t-1},t-1})$. This implies $\mathbb{P}(a_{t}=a_{\star}=\mathbb{P}(a_{t}=a^{\prime})$ and
$\mathbb{P}(a_{t}=a_{\star})=1-\mathbb{P}(a_{t}=a^{\prime})-\mathbb{P}(a_{t}=a_{o})\leq 1-\mathbb{P}(a_{t}=a^{\prime})=1-\mathbb{P}(a_{t}=a_{\star}),$
and the probability of selecting an optimal arm cannot exceed $1/2$. Thus, the expected regret,
$\mathrm{Reg}_{\Pi}(T,\theta_{\star},\mathbf{z}_{a_{\star}},\mathbf{z}_{a^{\prime}},\mathbf{z}_{a_{o}})\geq\Big{(}\frac{1}{2}-\frac{1}{6}\Big{)}\sum_{t=1}^{T}\mathbb{P}(a_{t}\neq a_{\star})\geq\frac{T}{6},$
which completes the proof.
Appendix E Technical Lemmas
Lemma 1.
(Exponential martingale inequality) If a martingale $(\mathbf{X}_{t};t\geq 0)$, adapted to filtration $\mathcal{F}_{t}$, satisfies $\mathbb{E}[{\exp(\lambda\mathbf{X}_{t})}{\hskip 0.80002pt|\hskip 0.80002pt}{\mathcal{F}_{t-1}}]\leq\exp(\lambda^{2}\sigma_{t}^{2}/2)$ for some constant $\sigma_{t}$, for all $t$, then for any $a\geq 0$,
$\mathbb{P}\left(|{X_{T}-X_{0}}|\geq a\right)\leq 2\exp\left(-\frac{a^{2}}{2\sum_{t=1}^{T}\sigma_{t}^{2}}\right)$
Thus, with probability at least $1-\delta$,
$|{X_{T}-X_{0}}|\leq\sqrt{2\sum_{t=1}^{T}\sigma_{t}^{2}\log\frac{2}{\delta}}.$
E.1 A Hoeffding bound for Matrices
Lemma 2.
[…]
$\mathbb{P}\left(\left\|\sum_{\tau=1}^{t}\mathbf{M}_{\tau}-\mathbb{E}[{\mathbf{M}_{\tau}}{\hskip 0.80002pt|\hskip 0.80002pt}{\mathcal{F}_{\tau-1}}]\right\|_{2}\geq x\right)\leq 2d\exp\left(-\frac{x^{2}}{2tb^{2}}\right)$
Proof.
[…]
$\mathbb{P}\left(\left\|\sum_{\tau=1}^{t}\mathbf{D}_{\tau}\right\|_{2}\geq x\right)\leq\mathbb{P}\left(\lambda_{\max}\left(\sum_{\tau=1}^{t}\mathbf{D}_{\tau}\right)\geq x\right)+\mathbb{P}\left(\lambda_{\max}\left(-\sum_{\tau=1}^{t}\mathbf{D}_{\tau}\right)\geq x\right)$
We bound the first term and the second term is bounded with similar arguement. For any $v>0$,
$\mathbb{P}\left(\lambda_{\max}\left(\sum_{\tau=1}^{t}\mathbf{D}_{\tau}\right)\geq x\right)\leq\mathbb{P}\left(\exp\left\{v\lambda_{\max}\left(\sum_{\tau=1}^{t}\mathbf{D}_{\tau}\right)\right\}\geq e^{vx}\right)\leq e^{-vx}\mathbb{E}\left[\exp\left\{v\lambda_{\max}\left(\sum_{\tau=1}^{t}\mathbf{D}_{\tau}\right)\right\}\right].$
Since $\sum_{\tau=1}^{t}\mathbf{D}_{\tau}$ is a real symmetric matrix,
$\exp\left\{v\lambda_{\max}\left(\sum_{\tau=1}^{t}\mathbf{D}_{\tau}\right)\right\}=$ $\lambda_{\max}\left\{\exp\left(v\sum_{\tau=1}^{t}\mathbf{D}_{\tau}\right)\right\}\leq\text{Tr}\left\{\exp\left(v\sum_{\tau=1}^{t}\mathbf{D}_{\tau}\right)\right\},$
where the last inequality holds since $\exp(v\sum_{\tau=1}^{t}\mathbf{D}_{\tau})$ has nonnegative eigenvalues. Taking expectation on both side gives,
$\mathbb{E}\left[\exp\left\{v\lambda_{\max}\left(\sum_{\tau=1}^{t}\mathbf{D}_{\tau}\right)\right\}\right]$ $\leq\mathbb{E}\left[\text{Tr}\left\{\exp\left(v\sum_{\tau=1}^{t}\mathbf{D}_{\tau}\right)\right\}\right]$
$=\text{Tr}\mathbb{E}\left[\exp\left(v\sum_{\tau=1}^{t}\mathbf{D}_{\tau}\right)\right]$
$=\text{Tr}\mathbb{E}\left[\exp\left(v\sum_{\tau=1}^{t-1}\mathbf{D}_{\tau}+\log\exp(v\mathbf{D}_{t})\right)\right].$
By Lieb’s theorem [Tropp, 2015] the mapping $\mathbf{D}\mapsto\exp(\mathbf{H}+\log\mathbf{D})$ is concave on positive symmetric matrices for any symmetric positive definite $H$. By Jensen’s inequality,
$\text{Tr}\mathbb{E}\left[\exp\left(v\sum_{\tau=1}^{t-1}\mathbf{D}_{\tau}+\log\exp(v\mathbf{D}_{t})\right)\right]\leq\text{Tr}\mathbb{E}\left[\exp\left(v\sum_{\tau=1}^{t-1}\mathbf{D}_{\tau}+\log\mathbb{E}\left[\left.\exp(v\mathbf{D}_{t})\right|\mathcal{F}_{t-1}\right]\right)\right]$
By Hoeffding’s lemma,
$e^{vx}\leq\frac{b-x}{2b}e^{-vb}+\frac{x+b}{2b}e^{vb}$
for all $x\in[-b,b]$. Because the eigenvalue of $\mathbf{D}_{\tau}$ lies in $[-b,b]$, we have
$\mathbb{E}\left[\left.\exp(v\mathbf{D}_{t})\right|\mathcal{F}_{t-1}\right]$ $\preceq\mathbb{E}\left[\left.\frac{e^{-vb}}{2b}\left(b\mathbf{I}_{d}-\mathbf{D}_{t}\right)+\frac{e^{vb}}{2b}\left(\mathbf{D}_{t}+b\mathbf{I}_{d}\right)\right|\mathcal{F}_{t-1}\right]$
$=\frac{e^{-vb}+e^{vb}}{2}\mathbf{I}_{d}$
$\preceq\exp(\frac{v^{2}b^{2}}{2})\mathbf{I}_{d}.$
Recursively,
$\mathbb{E}\left[\exp\left\{v\lambda_{\max}\left(\sum_{\tau=1}^{t}\mathbf{D}_{\tau}\right)\right\}\right]$ $\leq\text{Tr}\mathbb{E}\left[\exp\left(v\sum_{\tau=1}^{t-1}\mathbf{D}_{\tau}+\log\mathbb{E}\left[\left.\exp(v\mathbf{D}_{t})\right|\mathcal{F}_{t-1}\right]\right)\right]$
$\leq\text{Tr}\mathbb{E}\left[\exp\left(v\sum_{\tau=1}^{t-1}\mathbf{D}_{\tau}+(\frac{v^{2}b^{2}}{2})\mathbf{I}_{d}\right)\right]$
$\leq\text{Tr}\mathbb{E}\left[\exp\left(v\sum_{\tau=1}^{t-2}\mathbf{D}_{\tau}+(\frac{v^{2}b^{2}}{2})\mathbf{I}_{d}+\log\mathbb{E}\left[\left.\exp(vD_{t-1})\right|\mathcal{F}_{t-2}\right]\right)\right]$
$\leq\text{Tr}\mathbb{E}\left[\exp\left(v\sum_{\tau=1}^{t-2}\mathbf{D}_{\tau}+(\frac{2v^{2}b^{2}}{2})\mathbf{I}_{d}\right)\right]$
[…]
$\leq\text{Tr}\exp\left((\frac{tv^{2}b^{2}}{2})\mathbf{I}_{d}\right)$
$=\exp\left(\frac{tv^{2}b^{2}}{2}\right)\text{Tr}\left(\mathbf{I}_{d}\right)$
$=d\exp\left(\frac{tv^{2}b^{2}}{2}\right).$
Thus we have
$\mathbb{P}\left(\lambda_{\max}\left(\sum_{\tau=1}^{t}\mathbf{D}_{\tau}\right)\geq x\right)\leq d\exp\left(-vx+\frac{tv^{2}b^{2}}{2}\right).$
Minimizing over $v>0$ gives $v=x/(tb^{2})$ and
$\mathbb{P}\left(\lambda_{\max}\left(\sum_{\tau=1}^{t}\mathbf{D}_{\tau}\right)\geq x\right)\leq d\exp\left(-\frac{x^{2}}{2tb^{2}}\right),$
which proves the lemma. ∎
E.2 A Bound for the Gram Matrix
The Hoeffding bound for matrices (Lemma 2) implies the following bound for the two Gram matrices $\mathbf{A}_{t}:=\sum_{\tau=1}^{t}\widetilde{\mathbf{x}}_{a_{\tau}}\widetilde{\mathbf{x}}_{a_{\tau}}^{\top}$ and $\mathbf{V}_{t}:=\sum_{\tau=1}^{t}\sum_{a\in[K]}\widetilde{\mathbf{x}}_{a}\widetilde{\mathbf{x}}_{a}^{\top}$
Corollary 1.
For any $\epsilon\in(0,1)$ and $t\geq 8\epsilon^{-2}(1-p)^{-2}K^{2}\log\frac{2Kt^{2}}{\delta}$, with probability at least $1-\delta/t^{2},$
$\left\|\mathbf{I}_{K}-\mathbf{V}_{t}^{-1/2}\mathbf{A}_{t}\mathbf{V}_{t}^{-1/2}\right\|_{2}\leq\epsilon,$
Proof.
Note that
$\mathbf{V}_{t}^{-1/2}\mathbf{A}_{t}\mathbf{V}_{t}^{-1/2}-\mathbf{I}_{K}=\mathbf{V}_{t}^{-1/2}\left\{\sum_{\tau=1}^{t}\sum_{a\in[K]}\left(\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,\tau}}-1\right)\mathbf{\widetilde{x}}_{a}\mathbf{\widetilde{x}}_{a}^{\top}\right\}\mathbf{V}_{t}^{-1/2},$
and the martingale difference matrix for each $\tau\in[t]$,
$\left\|\sum_{a\in[K]}\left(\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,\tau}}-1\right)\mathbf{V}_{t}^{-1/2}\mathbf{\widetilde{x}}_{a}\mathbf{\widetilde{x}}_{a}^{\top}\mathbf{V}_{t}^{-1/2}\right\|_{2}$ $\leq\left(\frac{K-1}{1-p}+K-2\right)\max_{a\in[K]}\left\|\mathbf{V}_{t}^{-1/2}\mathbf{\widetilde{x}}_{a}\mathbf{\widetilde{x}}_{a}^{\top}\mathbf{V}_{t}^{-1/2}\right\|_{2}$
$\leq\frac{2K}{1-p}\max_{a\in[K]}\left\|\mathbf{\widetilde{x}}_{a}\right\|_{\mathbf{V}_{t}^{-1}}^{2}$
$\leq\frac{2K}{1-p}\cdot\frac{1}{t},$
where the last inequality holds by Sherman-Morrison formula. By Hoeffding bound for matrix (Lemma 2), for $x>0$
$\mathbb{P}\left(\left\|\mathbf{V}_{t}^{-1/2}\mathbf{A}_{t}\mathbf{V}_{t}^{-1/2}-\mathbf{I}_{K}\right\|_{2}>x\right)\leq 2K\exp\left(-\frac{(1-p)^{2}tx^{2}}{8K^{2}}\right).$
Setting $x=\epsilon\in(0,1)$ which will be determined later, for $t\geq 8\epsilon^{-2}(1-p)^{-2}K^{2}\log\frac{2Kt^{2}}{\delta}$ with probability at least $1-\delta/t^{2},$
$\left\|\mathbf{I}_{K}-\mathbf{V}_{t}^{-1/2}\mathbf{A}_{t}\mathbf{V}_{t}^{-1/2}\right\|_{2}\leq\epsilon,$
E.3 An error bound for the Lasso estimator
Lemma 3 (An error bound for the Lasso estimator with unrestricted minimum eigenvalue).
Let $\{\mathbf{x}_{\tau}\}_{\tau\in[t]}$ denote the covariates in $[-1,1]^{d}$ and $y_{\tau}=\mathbf{x}_{\tau}^{\top}\bar{\mathbf{w}}+e_{\tau}$ for some $\bar{\mathbf{w}}\in\mathbb{R}^{d}$ and $e_{\tau}\in\mathbb{R}$. For $\lambda>0$, let
$\widehat{\mathbf{w}}_{t}=\mathop{\mathrm{argmin}}_{\mathbf{w}}\sum_{\tau=1}^{t}\left(y_{\tau}-\mathbf{x}_{\tau}^{\top}\mathbf{w}\right)^{2}+\lambda\|\mathbf{w}\|_{1}.$
Let $\bar{\mathcal{S}}:=\{i\in[d]:\bar{\mathbf{w}}(i)\neq 0\}$ and $\boldsymbol{\Sigma}_{t}:=\sum_{\tau=1}^{t}\mathbf{x}_{\tau}\mathbf{x}_{\tau}^{\top}$. Suppose $\boldsymbol{\Sigma}_{t}$ has positive minimum eigenvalue and $\|\sum_{\tau=1}^{t}e_{\tau}\mathbf{x}_{\tau}\|_{\infty}\leq\lambda/2$. Then,
$\|\widehat{\mathbf{w}}_{t}-\bar{\mathbf{w}}\|_{\boldsymbol{\Sigma}_{t}}\leq\frac{2\lambda\sqrt{|{\bar{\mathcal{S}}}|}}{\sqrt{\lambda_{\min}\left(\boldsymbol{\Sigma}_{t}\right)}}.$
Proof.
The proof is similar to that of Lemma B.4 in [Kim et al., 2024], but we provide a new proof for the (unrestricted) minimum eigenvalue condition. Let $\mathbf{X}_{t}^{\top}:=(\mathbf{x}_{1},\ldots,\mathbf{x}_{t})\in[-1,1]^{d\times t}$ and $\mathbf{e}_{t}^{\top}:=(e_{1},\ldots,e_{t})\in\mathbb{R}^{t}$. We write $\mathbf{X}_{t}(j)$ and $\widehat{\mathbf{w}}_{t}(j)$ as the $j$-th column of $\mathbf{X}_{t}$ and $j$-th entry of $\widehat{\mathbf{w}}_{t}$, respectively. By definition of $\widehat{\mathbf{w}}_{t}$,
$\|\mathbf{X}_{t}\left(\bar{\mathbf{w}}-\widehat{\mathbf{w}}_{t}\right)+\mathbf{e}_{t}\|_{2}^{2}+\lambda\|\widehat{\mathbf{w}}_{t}\|_{1}\leq\|\mathbf{e}_{t}^{(j)}\|_{2}^{2}+\lambda\|\bar{\mathbf{w}}\|_{1},$
which implies
$\|\mathbf{X}_{t}\left(\bar{\mathbf{w}}-\widehat{\mathbf{w}}_{t}\right)\|_{2}^{2}+\lambda\|\widehat{\mathbf{w}}_{t}\|_{1}$ $\leq 2\left(\widehat{\mathbf{w}}_{t}-\bar{\mathbf{w}}\right)^{\top}\mathbf{X}_{t}^{\top}\mathbf{e}_{t}+\lambda\|\bar{\mathbf{w}}\|_{1}$
$\leq 2\|\widehat{\mathbf{w}}_{t}-\bar{\mathbf{w}}\|_{1}\|\mathbf{X}_{t}^{\top}\mathbf{e}_{t}\|_{\infty}+\lambda\|\bar{\mathbf{w}}\|_{1}$
$\leq\lambda\|\widehat{\mathbf{w}}_{t}-\bar{\mathbf{w}}\|_{1}+\lambda\|\bar{\mathbf{w}}\|_{1},$
where the last inequality uses the bound on $\lambda$. On the left hand side, by triangle inequality,
$\|\widehat{\mathbf{w}}_{t}\|_{1}=$ $\sum_{i\in\bar{\mathcal{S}}}|{\widehat{\mathbf{w}}_{t}(i)}|+\sum_{i\in[d]\setminus\bar{\mathcal{S}}}|{\widehat{\mathbf{w}}_{t}(i)}|$
$\geq$ $\sum_{i\in\bar{\mathcal{S}}}|{\widehat{\mathbf{w}}_{t}(i)}|-\sum_{i\in\mathcal{S}_{\star}}|{\widehat{\mathbf{w}}_{t}(i)-\bar{\mathbf{w}}(i)}|+\sum_{i\in[d]\setminus\bar{\mathcal{S}}}|{\bar{\mathbf{w}}(i)}|$
$=$ $\|\bar{\mathbf{w}}\|_{1}-\sum_{i\in\bar{\mathcal{S}}}|{\widehat{\mathbf{w}}_{t}(i)-\bar{\mathbf{w}}(i)}|+\sum_{i\in[d]\setminus\bar{\mathcal{S}}}|{\widehat{\mathbf{w}}_{t}(i)}|$
and for the right-hand side,
$\|\widehat{\mathbf{w}}_{t}-\bar{\mathbf{w}}\|_{1}=\sum_{i\in\bar{\mathcal{S}}}|{\widehat{\mathbf{w}}_{t}(i)-\bar{\mathbf{w}}(i)}|+\sum_{i\in[d]\setminus\bar{\mathcal{S}}}|{\widehat{\mathbf{w}}_{t}(i)}|.$
Plugging in both sides and rearranging the terms,
$\|\mathbf{X}_{t}\left(\bar{\mathbf{w}}-\widehat{\mathbf{w}}_{t}\right)\|_{2}^{2}\leq 2\lambda\sum_{i\in\bar{\mathcal{S}}}|{\widehat{\mathbf{w}}_{t}(i)-\bar{\mathbf{w}}(i)}|.$ (26)
Because $\mathbf{X}_{t}^{\top}\mathbf{X}_{t}$ is positive definite,
$\|\mathbf{X}_{t}\left(\bar{\mathbf{w}}-\widehat{\mathbf{w}}_{t}\right)\|_{2}^{2}\geq$ $\lambda_{\min}(\mathbf{X}_{t}^{\top}\mathbf{X}_{t})\sum_{i\in\bar{\mathcal{S}}}|{\widehat{\mathbf{w}}_{t}(i)-\bar{\mathbf{w}}(i)}|^{2}$
$\geq$ $\frac{\lambda_{\min}(\mathbf{X}_{t}^{\top}\mathbf{X}_{t})}{|{\bar{\mathcal{S}}}|}\left(\sum_{i\in\bar{\mathcal{S}}}|{\widehat{\mathbf{w}}_{t}(i)-\bar{\mathbf{w}}(i)}|\right)^{2},$
where the last inequality holds by Cauchy-Schwarz inequality. Plugging in Eq. 26 gives,
$\|\mathbf{X}_{t}\left(\bar{\mathbf{w}}-\widehat{\mathbf{w}}_{t}\right)\|_{2}^{2}\leq$ $2\lambda\sum_{i\in\bar{\mathcal{S}}}|{\widehat{\mathbf{w}}_{t}(i)-\bar{\mathbf{w}}(i)}|$
$\leq$ $2\lambda\sqrt{\frac{|{\bar{\mathcal{S}}}|}{\lambda_{\min}(\boldsymbol{\Sigma}_{t})}}\|\mathbf{X}_{t}\left(\bar{\mathbf{w}}-\widehat{\mathbf{w}}_{t}\right)\|_{2}$
$\leq$ $\frac{2\lambda^{2}|{\bar{\mathcal{S}}}|}{\lambda_{\min}(\boldsymbol{\Sigma}_{t})}+\frac{1}{2}\|\mathbf{X}_{t}\left(\bar{\mathbf{w}}-\widehat{\mathbf{w}}_{t}\right)\|_{2}^{2},$
where the last inequality uses $ab\leq a^{2}/2+b^{2}/2$. Rearranging the terms,
$\|\mathbf{X}_{t}\left(\bar{\mathbf{w}}-\widehat{\mathbf{w}}_{t}\right)\|_{2}^{2}\leq\frac{4\lambda^{2}|{\bar{\mathcal{S}}}|}{\lambda_{\min}(\boldsymbol{\Sigma}_{t})},$
which proves the result. ∎
E.4 Eigenvalue bounds for the Gram matrix.
Lemma 4.
For $a\in[K]$, let $\widetilde{\mathbf{x}}_{a}:=[\mathbf{x}_{a}^{\top},\mathbf{e}_{a}^{\top}\mathbf{p}_{1},\cdots,\mathbf{e}_{a}^{\top}\mathbf{p}_{K-d}]^{\top}\in\mathbb{R}^{d}$ denote augmented features. Then, an eigenvalue of $\sum_{a\in[K]}\widetilde{\mathbf{x}}_{a}\widetilde{\mathbf{x}}_{a}^{\top}$ is in the following interval
$\left[\min\left\{\lambda_{\min}\left(\sum_{a\in[k]}\mathbf{x}_{a}\mathbf{x}_{a}^{\top}\right),1\right\},\max\left\{\lambda_{\max}\left(\sum_{a\in[K]}\mathbf{x}_{a}\mathbf{x}_{a}^{\top}\right),1\right\}\right].$
Proof.
Let $\mathbf{P}:=(\mathbf{p}_{1},\ldots,\mathbf{p}_{K-d})\in\mathbb{R}^{K\times(K-d)}$. Because the columns in $\mathbf{P}$ are orthogonal each other and to $\mathbf{x}_{1},\ldots,\mathbf{x}_{K}$,
$\sum_{a\in[K]}\widetilde{\mathbf{x}}_{a}\widetilde{\mathbf{x}}_{a}^{\top}=$ $\begin{bmatrix}\sum_{a\in[K]}\mathbf{x}_{a}\mathbf{x}_{a}^{\top}&\sum_{a\in[K]}\mathbf{x}_{a}\mathbf{e}_{a}^{\top}\mathbf{P}\\ \sum_{a\in[K]}\mathbf{P}^{\top}\mathbf{e}_{a}\mathbf{x}_{a}^{\top}&\mathbf{P}^{\top}\mathbf{P}\end{bmatrix}$
$=$ $\begin{bmatrix}\sum_{a\in[K]}\mathbf{x}_{a}\mathbf{x}_{a}^{\top}&\sum_{a\in[K]}\mathbf{x}_{a}\mathbf{e}_{a}^{\top}\mathbf{P}\\ \sum_{a\in[K]}\mathbf{P}^{\top}\mathbf{e}_{a}\mathbf{x}_{a}^{\top}&I_{K-d}\end{bmatrix}$
$=$ $\begin{bmatrix}\sum_{a\in[K]}\mathbf{x}_{a}\mathbf{x}_{a}^{\top}&\sum_{a\in[K]}\mathbf{X}\mathbf{e}_{a}\mathbf{e}_{a}^{\top}\mathbf{P}\\ \sum_{a\in[K]}\mathbf{P}^{\top}\mathbf{e}_{a}\mathbf{e}_{a}^{\top}\mathbf{X}&I_{K-d}\end{bmatrix}$
$=$ $\begin{bmatrix}\sum_{a\in[K]}\mathbf{x}_{a}\mathbf{x}_{a}^{\top}&\mathbf{X}\mathbf{P}\\ \mathbf{P}^{\top}\mathbf{X}^{\top}&I_{K-d}\end{bmatrix}$
$=$ $\begin{bmatrix}\sum_{a\in[K]}\mathbf{x}_{a}\mathbf{x}_{a}^{\top}&\mathbf{O}\\ \mathbf{O}&I_{K-d}\end{bmatrix}.$
Thus, for any $\lambda\in\mathbb{R}$, $\det(\sum_{a\in[K]}\widetilde{\mathbf{x}}_{a}\widetilde{\mathbf{x}}_{a}^{\top}-\lambda\mathbf{I}_{K})=\det(\sum_{a\in[K]}\mathbf{x}_{a}\mathbf{x}_{a}^{\top}-\lambda\mathbf{I}_{d})(1-\lambda)^{K-d}$. Solving $\det(\sum_{a\in[K]}\mathbf{x}_{a}\mathbf{x}_{a}^{\top}-\lambda\mathbf{I}_{d})(1-\lambda)^{K-d}=0$ gives the eigenvalues and the lemma is proved. ∎